### v2

# Persuasion in Networks: Can the Sender Do Better than Using Public Signals?

Abstract
Political and advertising campaigns increasingly exploit social networks to spread information and persuade people. This paper studies a persuasion model to examine whether such a strategy is better than simply sending public signals. Receivers in the model have heterogeneous priors and will pass on a signal if they are persuaded by it to take sender’s preferred action. I show that a risk neutral or risk loving sender prefers to use public signals, unless more skeptical receivers are sufficiently more connected in the social network. A risk averse sender may prefer to exploit the network. These results still hold when the network exhibits homophily.
Keywords: Bayesian persuasion, networks, random networks, diffusion
People are increasingly relying on social networks to gather information. Many argue that information going viral on social media had significant influences on people’s decisions related to elections and public health [^2] See e.g. Bogart and Thorburn (2005), Allcott and Gentzkow (2017) and references therein . Realising the power of such viral information, firms and politicians have spent a vast amount of resources on social media campaigns. Their spending on social media is also increasing relative to traditional mass media [^3] See, for example, Nelson et al. (2021) for political advertisements and the 28th edition of CMO survey for commercial advertisements( https://cmosurvey.org/results/28th-edition-february-2022/ ) . A natural question is how effective such social media campaigns are in influencing people’s decisions. In particular, are they more effective than mass media?
To see the trade-off between using mass media and social networks, consider an example where a pharmaceutical firm tries to advertise its new drug. Some customers are skeptical about the drug while others have prior belief that the drug is safe and effective. The firm would like to send the result of a clinical trial to the skeptics, but keep such information away from the believers, since negative results from the trials could dissuade the believers. If mass media is used to disseminate information, everyone must receive the same information. With information disseminated on social networks, the firm could try to target information towards the sceptics. However, people will also be sharing information with each other, so it is unclear whether and when the firm can succeed in sending information more often to sceptics than believers, and thus unclear whether the firm prefers public dissemination or dissemination using social network. In addition, the risk properties of the two options are different. Everyone with the same prior must choose the same action with public dissemination, whereas people can have different observations and thus use different actions when information is spread on a network.
To address these questions, I propose the following persuasion model. There are two states and receivers have two actions $\{0,1\}$. There are infinitely many receivers, who want to match their action with the state. The sender wants receivers to take action 1 regardless of the state. Receivers have heterogeneous priors, some of them (believers) would choose action 1 while others (sceptics) would choose action 0 if no further information is provided. Receivers are located on a network captured by a random network model. Receivers know their own degree.
The sender can use public signals or network signals. With public signals, all receivers must receive the same realised signal. If the sender uses network signals, then the sender can select a small proportion of receivers as seeds based on their priors and positions in the network. These seeds will receive signals from the sender. When receivers observe some signal, they decide whether to pass it on to their neighbours. In the main analysis, I assume that receivers pass on a signal if and only if they are persuaded by it to take the sender’s preferred action. Intuitively, if people are persuaded to purchase a product or vote for a politician, they are willing to share its advertisement to their friends.
I allow the sender to send no signal. This plays no role with public signals. If sender uses network signals, receivers may receive no information due to the network structure.When receivers observe no signal, they cannot distinguish whether it is due to network structure or sender choosing to send no signal. The sender could potentially exploit this in her benefit.
The main question in this paper is whether and when network signals can outperform public signals for the sender. The potential cost of network signals is that some sceptics may not receive information due to structure of the network and/or how signals are shared. The potential benefit of network signal is that the sender may be able to send information more often to the sceptics than to believers. Also, as noted above, network and public signals have different risks. Conditions under which the potential benefit can and cannot be achieved are identified. This in turn allows us to compare public and network signals, taking into account sender’s risk preferences.
The first main result states that in the baseline network (i.e. one without homophily or correlations between degree and priors), if the sender is risk loving or risk neutral, network signals cannot outperform public signals. It is possible for network signals to perform strictly better if the sender is risk averse. Believers and skeptics will be ex ante symmetric in terms of their positions in such networks. Therefore, there are no differences between how skeptics and believers are exposed to signals, and the sender cannot send information more often to sceptics than believers. The sender can only prefer network signals due to its risk properties.
Network signals in general result in receivers sometimes not observing any signal. With public signals, if receivers do not receive good news, they know with certainty that good news is not realised. With network signals, the event of not observing good news is less informative since it may be due to the network structure. This allows the sender to increase the probability of sending good news, without dissuading believers when good news is not realised. This comes at the cost that when good news does realise, less receivers observe it. Therefore, network signals allow the sender to reduce risk, so could be beneficial if she is risk averse.
Introducing homophily or making believers more connected than sceptics does not change these results. Networks exhibiting these features do introduce asymmetry between believers and sceptics. However, the asymmetry will actually make believers more exposed to signals than sceptics, which is the opposite of what the sender desires.
If skeptics are sufficiently more connected in the network than believers, the sender prefers network signals regardless of her risk preferences. Skeptics being more connected allows the sender to send information to them more often than to the believers.
In summary, this paper compares performance of public and network signals. In particular, it shows that under many plausible conditions, public signals perform better. Network signals can be preferred by the sender only when the sender is risk averse and/or sceptics are more connected. In practice, sender in many situations have access to tools for public dissemination of information. The results imply that, in many situations, resources may be more efficiently spent on sending public signals, rather than on trying to understand network structures and picking the optimal seeds. In addition, persuasion on large networks is complicated while public signals have been well studied. From these comparisons, we can use sender’s value from public signals to obtain what sender could get at most or at least using network signals, whose value for the sender can be very difficult to compute.
Relation and contribution to the literature are discussed in the next section. Section 2 presents the model. Section 4 presents the main results, and Section 3 contains lemmas useful for understanding the main results. Section 5 discusses implications of relaxing some assumptions and possible extensions. The final section concludes.
1 Literature Review
This paper is most related to the literature on persuasion with information spillovers. Galperti and Perego (2023) studies an environment where a sender can directly send signals to some receivers (seeds), and receivers observe the same information as those connected to them (by a directed path). They show that the problem is equivalent to allowing the sender to directly send information to all receivers, by transforming the network appropriately. They provide a characterization of feasible outcomes and show how it changes with the network and seeds. Babichenko et al. (2021) studies a similar problem, and also explores the algorithmic approach. Kerman and Tenev (2021) studies a voting game in which receivers observe their neighbours’ signals. They characterise sender’s optimal signals in some stylised networks, and compare sender’s payoff to the case when receivers do not communicate. Liporace (2021) assumes that receivers observe signals of their neighbours, and receivers have heterogeneous priors. She characerises sender’s optimal strategy, restricting attention to strategies that persuade everyone in the good state. In Egorov and Sonin (2020), information is shared between receivers with some probability, and receivers choose whether to subscribe to the sender which is costly. They characterise sender’s optimal strategy in some stylised networks.
This strand of literature usually assumes that information will be transmitted mechanically to one’s neighbours, while this paper studies more endogenous spread of information. Another key difference is that, the sender in these papers can always make sure that all receivers observe information from the sender, so they do not consider the possibility that sender’s information may not reach some receivers, which is very likely when considering large social networks. Also, in this paper, sender and receivers do not have full information of the network, whereas the literature studies fixed and known networks. Uncertainties over network is a plausible assumption in large networks.
Also related is the literature on strategic communication, which studies how players communicate when their payoffs depend on each other’s actions. Most papers in this literature (e.g. Galeotti et al. (2013) and Hagenbach and Koessler (2010)) focus on conditions under which players truthfully communicate with each other, and do not study persuasion problems. The exception is Squintani (2024). Squintani (2024) studies an environment where there is one sender and one receiver (randomly selected from all players). He characterises conditions for truthful communication and studies the optimal network and network formation.
Some papers have looked at the comparison between public signals and private signals, but do not allow communication between receivers. Wang (2013) compares public signals and i.i.d. signals in a voting game, and finds that public signal is always better. Inostroza and Pavan (2018) also compares public and private signals, imposing certain restrictions on private signals. Candogan (2020) studies a problem of a social media platform to signal the quality of content, and finds that public signals are sufficient if the goal is to minimize misinformation. Arieli and Babichenko (2019) relates optimality of public signals to the core of a cooperative game. Communication between receivers puts endogenous restrictions on joint distributions over what receivers observe when sender sends personalised signals. Also, receivers have heterogeneous priors in this paper. This potentially makes personalised signals more attractive. Due to these and other more detailed differnces in setup, the trade-offs and insights in this paper are different from existing papers in this literature.
Finally, this paper uses random networks to model the social network. Random networks have been used in Akbarpour et al. (2020), Sadler (2020), Campbell (2013) and Golub and Jackson (2012) among others, to study various problems.
2 Model
2.1 Agents and Network
I consider an environment with one sender, and an infinite and countable set of receivers. There are two possible states of the world, $\omega\in\{0,1\}$, and each receiver has two possible actions $\{0,1\}$. The receivers want to match their action with the state. The receiver gets a payoff of 1 if their action matches the state, and gets 0 otherwise. The sender wants receivers to choose action 1 regardless of the state. If a proportion $x$ of the receivers choose action 1, then the sender gets payoff $v(x)$, where $v(.)$ is continuous, weakly increasing and bounded. I normalise $v(0)=0$.
The receivers’ payoff function implies that action 1 is optimal for receivers if and only if they believe the probability of state 1 is greater than 0.5. Throughout the paper, I break indifference in favour of the sender [^4] Results can be extended if I break indifference in the other way, and proofs are essentially the same., i.e. I assume that receivers choose action 1 when they believe the probability of state 1 is exactly 0.5.
A proportion $\gamma_{h}$ of the receivers are of type $h$ and have prior belief that state $\omega$ realises with probability $\mu_{h}(\omega)$. The rest of the receivers are of type $l$, and have a prior $\mu_{l}(\omega)$ for state $\omega$. I assume that $\mu_{h}(\omega=1)>0.5>\mu_{l}(\omega=1)$. I denote the proportion of type $l$ receivers by $\gamma_{l}$. Type $h$ receivers are often referred to as believers, and type $l$ receivers as sceptics.
The sender has a prior belief that state $\omega$ realises with probability $\mu_{s}(\omega)$. The sender can send signals from the signal space $S\cup\{\emptyset\}$, where $S$ is finite and contains at least two elements [^5] All results below easily extend to cases where $S$ has only one element or $S$ is empty. If $S$ is empty, the result is trivial since the sender essentially cannot do anything. If $S$ contains one element, payoff from network signals weakly decreases, while it can be easily shown that payoff using public signals does not change. For results where network signals perform better than public signals, in their proofs, sender only uses 1 signal from $S$ with positive probability.. The element $\emptyset$ captures the sender’s ability to send no signal. The sender designs a signal structure $\pi(.)$. The probability of signal $s\in S\cup\{\emptyset\}$ being realised in state $\omega$ is $\pi(s|\omega)$, and $\pi(.)$ must satisfy $\sum_{s\in S\cup\{\emptyset\}}\pi(s|\omega)=1$ for $\omega=0,1$. If the sender uses network signals, the sender can also choose where to seed the signals. Precisely how seeding works and timing of the game are described in the next two subsections.
Now we describe the random network model, which captures how receivers are connected to each other. Each receiver $i$ has connectedness $\lambda_{i}$. With $n$ receivers, if $i$ and $j$ are have the same prior, the probability that they are connected is $min\{\frac{\lambda_{i}\lambda_{j}}{n},1\}$. If they have different priors, the probability that they are connected is $min\{\frac{q\lambda_{i}\lambda_{j}}{n},1\}$ where $q\in(0,1]$. Intuitively, $q$ controls homophily. Whether $i$ and $j$ are connected is independent of connections between all other pairs (conditional on their connectedness and types). Conditional on having type $t$, a receiver has connectedness $\lambda$ with probability $f_{t}(\lambda)$, where $f_{t}(\cdot)$ is a probability mass function with finite support [^6] The analysis can be extended to the case where distribution over $\lambda$ is continuous and unbounded. This will make restrictions on these distributions and the proofs more tedious, without adding much insight. Bollobás et al. (2007) states the precise restrictions on these distributions when they have continuous and unbounded support..
Note that the probability that two nodes are connected depends on the number of receivers $n$. There are infinitely many receivers in the model, so we look at the limiting network when $n$ goes to infinity, as in many other papers using random network models. The network formation process is common knowledge. Each receiver $i$ in addition knows her own degree, denoted by $d_{i}$ (degree is the number of neighbours a receiver has). When studying network signals, I use receivers and nodes exchangeably.
Two important network statistics are degree distribution and the probability that a node is connected to a node of the same type. Theorem 3.13 in Bollobás et al. (2007) shows that
Remark 1.
As $n\rightarrow\infty$, degree distribution of type $t$ receivers converges to $\{p_{d}^{t}\}_{d}$. Also, conditional on degree $d$ and type $t$ of a node, each of its neighbour has an independent probability $q_{t}$ of having the same type. The degree distribution $\{p_{d}^{t}\}_{d}$ and probability $q_{t}$ are
$p_{d}^{t}=\sum_{\lambda^{\prime}}f_{t}(\lambda^{\prime})Poisson(d;\lambda^{\prime}(E_{t}(\lambda)\gamma_{t}+E_{t^{\prime}}(\lambda)\gamma_{t^{\prime}}q))\qquad q_{t}=\frac{E_{t}(\lambda)\gamma_{t}}{E_{t}(\lambda)\gamma_{t}+E_{t^{\prime}}(\lambda)\gamma_{t^{\prime}}q}$
where $E_{t}(\lambda)=\sum_{\lambda}f_{t}(\lambda)\lambda$ and $t^{\prime}\neq t$
2.2 Network signals
As described above, social network with infinitely many receivers is modelled as the limit of a sequence of finite random networks. Therefore, I will first describe the game with $n$ receivers. Sender’s optimal payoff facing infinitely many receivers will be captured as the limit of her optimal payoffs in these finite games.
With $n$ receivers, the timing goes as follows. The sender moves first and chooses a signal structure $\pi(.)$, and a seeding strategy $Z(.)$. Given a network with $n$ receivers $g^{n}$ and a signal $s$, $Z(g^{n},s)$ returns a subset of receivers. The subset selected by $Z(.)$ must have a smaller size than $n^{\alpha}$, where $\alpha<1$. Intuitively, $Z(.)$ selects some seeds that will receive signals first, and seed selection can be conditioned on the network and the signal realised.
After seeing the sender’s strategy, the receivers choose their actions for each degree $d$ and observation in $S\cup\{\emptyset\}$, $a_{i}:(S\cup\emptyset)\times\mathbb{N}\rightarrow\{0,1\}$, where $a_{i}(s,d)$ describes what $i$ would do when she has degree $d$ and observes $s\in S\cup\{\emptyset\}$.
Then nature draws a network $g^{n}$ according to the network formation process described above. Nature draws a state, a signal $s$ according to $\pi(.)$. If the signal is $\emptyset$, all receivers observe $\emptyset$. Intuitively, if the sender sends out no signal, then all receivers observe no signal.
If $s\neq\emptyset$, then seeds are selected according to $Z(.)$. Given network $g^{n}$ and the seeds, a receiver $i$ observes $s$ if and only if at least one of the following conditions hold:
  1. 1.
    $i$ is a seed.
  2. 2.
    $i$ has a neighbour $j$ who observes $s$, and $j$ chooses action 1 upon observing $s$,
    i.e. $a_{j}(s,d_{j})=1$.
All the other receivers observe $\emptyset$. In words, the condition here says that a receiver will pass on a signal to her neighbours, if and only if she is persuaded by the signal to take sender’s preferred action. Intuitively, people will pass on the advertisement of a product or a politician if they are convinced to buy the product or vote for the politician.
The solution concept is subgame perfect equilibrium. Note that, although we assume that the receivers move before realisation of their degree and signals, we are allowing their actions to be conditioned upon these information. Therefore, it is essentially the same as assuming that they move after observing their degrees and the signal. This is standard in Bayesian Persuasion (Kamenica and Gentzkow, 2011) and when studying Bayesian games more generally.
For each $n$, we can find the sender’s maximum (or supremum) equilibrium payoff. Denote this payoff by $V^{*}_{net}(n)$. Sender’s payoff facing infinitely many receivers is captured by $lim_{n\rightarrow\infty}V^{*}_{net}(n)$. However, this limit may not exist for some parameters [^7] It can be shown that this only happens for a non-generic set of parameters, but the proof is very long and does not bring additional insight, so I will not discuss it for the sake of space.. Therefore, I compare $limsup_{n\rightarrow\infty}V^{*}_{net}(n)$ and $liminf_{n\rightarrow\infty}V^{*}_{net}(n)$ with the sender’s payoff from public signals.
2.3 Public signals
Again, I describe the game with $n$ receivers, and sender’s optimal payoff facing infinitely many receivers is captured as the limit of her optimal payoffs in these finite games [^8] This is equivalent as just letting the sender face a continuum of receivers, of which $\gamma_{h}$ has type $h$. I describe sender’s problem as the sequence of finite games to be consistent with the setup above for network signals.
First, the sender designs and commits to signal structure $\pi(.)$. After seeing sender’s strategy, receivers choose their action upon observing each signal and $\emptyset$. This is captured by function $a_{i}:S\cup\emptyset\rightarrow\{0,1\}$.
The nature draws a state, and then a signal is realised according to $\pi(.)$. All receivers observe the realised signal. We study the subgame perfect equilibrium of this game. We can find sender’s optimal subgame perfect equilibrium payoff for each game with $n$ receivers, and I use $V^{*}_{pub}$ to denote the limit of sender’s optimal payoffs using public signals. The main research question of this paper is to compare $V^{*}_{pub}$ and limiting network signal payoffs defined in the previous subsection.
2.4 Discussion of Setup
Rule of sharing signals: The justification for the rule of sharing signals is that, after people purchase a certain product, they are more likely to talk about the product and thus pass information about it to their friends. Banerjee et al. (2013) studies diffusion of micro-finance and finds evidence that people who use micro-finance are more likely to transmit information about it to their neighbours, which provides indirect support for the rule used in this paper. Sharing behaviour used here is similar to those used in Campbell (2013) and Sadler (2020).
Also, main results are robust to other rules. In particular, another plausible rule of sharing is to share when the receivers’ action agrees with the meaning carried by the signal. This is discussed in more detail in section 5.
Seeds: I assume that seeds are a vanishing fraction of the population, mainly for tractability. This is nonetheless plausible on large social networks. Targeted advertising services on social media usually use auctions to select which advertisement to present to a user, and the auction takes into account the relevance of the advertisement to the user. Intending to target a wide audience using these services will result in losing more often to other competing ads, and it is difficult to actually reach a wide audience directly. Results can still go through if seeds as a fraction of population is non-vanishing but small enough. I discuss this and how results could change if the seeds is a large fraction of the population in section 5.
I am also allowing the sender to use information about the whole network when selecting seeds, and in practice they can use much less information. Results on network signals being worse than public signals are stronger if we restrict the information used in seeding. Results on network signals being better do not rely on sender’s ability to use so much information in seeding.
Empty Signal: From description of the public signal game, the empty signal’s role is not different from non-empty ones, so it is actually WLOG to assume that sender does not use empty signals in public signal game. For the network signal game, as we will see below, even if sender cannot use $\emptyset$, she can still essentially send no signal to almost everyone. Intuitively, this is done by choosing the seeds to be very isolated nodes. Allowing the sender to send $\emptyset$ directly is mainly for exposition. This also implies that one can think of $\emptyset$ as the sender choosing to conceal some signals. This interpretation requires sender to commit to concealment of signals before signal realisation. Relaxing the ability to commit to concealment plan restricts what sender can do using network signals, and thus strengthens results on public signals being better. For results on network signals being better, in the proofs for them, the sender uses strategies where non-empty signals give the sender a weakly higher payoff than $\emptyset$, so the sender has no incentive to conceal them.
3 Analysis of Receivers’ Posterior and the Network
In order to compare sender’s payoff from public and network signals, we first need to understand receivers’ best response facing sender’s strategies. This is straightforward for public signals, which is essentially the same as standard Bayesian persuasion. For network signals, analysis is more complicated since signals spread endogenously on networks. I present some lemmas that can be used to characterise limiting probabilities with which receivers observe each signal and $\emptyset$ when $n$ gets large. These lemmas can then be used to approximate and bound sender’s payoffs using network signals in large networks.
The next lemma pins down the receivers’ posterior upon observing some non-empty signal $s\neq\emptyset$. This lemma says if a receiver observes some non-empty signal $s$, then her posterior is the same as if all signals are sent publicly.
Lemma 1.
Given sender’s strategy $(\pi,Z)$ in game with $n$ receivers, posterior of a receiver with type $t$ and degree $d$ upon observing some signal $s\neq\emptyset$ is:
$\frac{\pi(s|\omega=1)\mu_{t}(\omega=1)}{\pi(s|\omega=1)\mu_{t}(\omega=1)+\pi(s|\omega=0)\mu_{t}(\omega=0)}$
The main idea of the proof is that network formation process and seeding strategy do not depend on the state, so the posterior upon observing some non-empty signal is determined completely by signal structure $\pi(.)$ and prior.
There are some important observations that follow from this lemma. If some signal $s$ persuades both types when sent publicly, then all receivers will share it if they receive it. If such a signal is realised, then a receiver will observe it if and only if she is connected to one of the seeds. In other words, all receivers on components [^9] A component in a network is a set of nodes such that all nodes in this set are connected to each other, and all nodes not in this set are not connected to nodes in the set. that contain at least one seed will observe the signal. If $s$ persuades only the believers but not the skeptics when sent publicly, then a receiver $i$ observes the signal if and only if there exists a path from $i$ to some seed such that all receivers on that path excluding $i$ are believers. In other words, all believers on components that contain at least one seed in the subnetwork of believers [^10] The subnetwork of believers is the network formed by deleting all sceptics and links from them will observe the signal, and a skeptic observes it if she has a neighbour who is a believer and observes the signal.
Receivers’ behaviour upon observing non-empty signals is directly implied by this lemma. We still need to understand receivers’ posterior upon observing $\emptyset$. To obtain these, we need the probabilities of observing $\emptyset$ in each state. I apply results from random networks to study probabilities of observing non-empty signals. The probability of observing $\emptyset$ can then be obtained as the complementary probability.
Let $L_{i}(g^{n})$ denote the $i$-th largest component in realised network $g^{n}$. Theorem 3.1 and 3.6 in Bollobás et al. (2007) show that
Theorem 1 (Bollobas et al. (2007)).
There exists $c\in[0,1]$ such that $|L_{1}(g^{n})|/n$ converges in probability to $c$, and $|L_{i}(g^{n})|/n$ converges in probability to 0 for $i>1$.
This theorem says that there is at most one component that is huge, in the sense that its size grows with $n$ linearly. The largest component $L_{1}$ is usually called the giant component.
As argued above, if a signal is only passed on by believers, we need to look at the subnetwork of believers. I use $\hat{L}_{i}(g^{n})$ to denote the $i$-th largest component in the believers’ subnetwork of $g^{n}$. The believers’ subnetwork depends only on connections between believers. Also, conditional on connectedness of a pair of believers, whether they are connected is independent of how sceptics are connected to each other and how sceptics are connected to believers. Therefore, the distribution over connections between believers, and thus the believers’ subnetwork, will not change if we set $\lambda_{i}$ to be 0 for all sceptics. Applying 1 then gives
Corollary 1.
There exists $\hat{c}\in[0,1]$ such that $|\hat{L}_{1}|/n$ converges in probability to $\hat{c}$, and $|\hat{L}_{i}|/n$ converges in probability to 0 for $i>1$.
Using these results on giant component and Lemma 1, we can specify some conditions under which the receivers observe $\emptyset$ with very high probability in large social networks.
Lemma 2.
Fix any $\epsilon>0$. For $n$ large enough, given any $s\in S$ and any seeding with no seeds on $L_{1}$, the proportion of receivers observing $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Also, if $s$ is only passed on by believers and no seeds are on $\hat{L}_{1}$, the proportion of receivers observing $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Intuitively, the lemma says that for a signal to reach many people, it must become viral, in the sense that it spreads on the giant component and thus is shared by a large fraction of the population. The main intuition of the proof is obvious from 1. Since all components other than the giant component contain a vanishing fraction of nodes, if no seeds hit the giant component, all components hit by the seeds have sizes that are vanishing as a proportion of the total population.
Lemma 3.
Fix any $\epsilon>0$. For $n$ large enough, given any $s\in S$ and any seeding, the proportion of receivers who are not on $L_{1}$ and observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Also, if $s$ is only passed on by believers, the proportion of receivers who have no neighbour on $\hat{L}_{1}$ and observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Intuitively, this lemma says that receivers can only observe the signal if they are on parts of the network where the signal has gone viral, i.e. the giant component. The main idea of proof is again intuitive. If some receiver is not on the giant component, then its component has vanishing size and is very unlikely to be hit by any seed.
Combining Lemma 3 and Lemma 2, the probability that a receiver can observe $s$ is approximately the probability that $s$ hits the appropriate giant component, times the probability that the receiver is on (or has a neighbour on) that giant component. The next lemma describes probabilities of being connected to $L_{1}$ and $\hat{L}_{1}$, conditional on type and degree of a node.
Lemma 4.
There exist functions $\zeta(t,d)$ and $\hat{\zeta}(t,d)$ both increasing in $d$ such that, conditional on having degree $d$ and type $t$:
  1. 1.
    the proportion of receivers on $L_{1}$ converges in probability to $\zeta(t,d)$.
  2. 2.
    the proportion of receivers with a neighbour on $\hat{L}_{1}$ converges in probability to $\hat{\zeta}(t,d)$.
The equations which characterise $\zeta(\cdot)$ and $\hat{\zeta}(\cdot)$ are not key to the results below, so are only presented in the appendix.
4 Results
4.1 Baseline network
In this subsection, I study the simplest case where the network formation process ignores the types of receivers. Formally, I study the case where $f_{t}(\cdot)$ is independent of $t$ and $q=1$.
The next lemma provides the main intuition underlying this subsection.
Lemma 5.
If $f_{l}(\lambda)=f_{h}(\lambda)$ for all $d$ and $q=1$, then we have $\zeta(h,d)=\zeta(l,d)$ and $\hat{\zeta}(h,d)=\hat{\zeta}(l,d)$ for all $d$.
This follows from the fact that, in this baseline network, the network formation process essentially ignores types of receivers. Intuitively, this says that if a signal gets viral, then believers and skeptics have the same probability of receiving it when $n$ is large. Therefore, when $n$ gets large, the probability of observing a signal is not different between the two types of receivers.
Proposition 1.
If $f_{l}(d)=f_{h}(d)$ for all $d$ and $q=1$, then
  1. 1.
    If $v(.)$ is linear or convex, $\limsup_{n\rightarrow\infty}V^{*}_{net}(n)\leq V^{*}_{pub}$.
  2. 2.
    If $\hat{c}>0$ (i.e. the giant component in type $h$ subnetwork is non-vanishing), there exist concave $v(.)$ such that $\liminf_{n\rightarrow\infty}V^{*}_{net}(n)>V^{*}_{pub}$.
For the second point, the following remark provides such a concave function $v(\cdot)$. It intuitively says that the sender prefers network signals if her risk aversion is high enough.
Remark 2.
Suppose that $v(x)=(x^{\beta}-1)/\beta$, which has constant relative risk aversion. If $\hat{c}>0$, the sender prefers network signals for sufficiently small $\beta$.
A necessary and sufficient condition for $\hat{c}>0$ can be found in Bollobás et al. (2007) and Akbarpour et al. (2020). A sufficient condition is that $\lambda_{min}^{2}\gamma_{h}\sum_{\lambda>0}f_{h}(\lambda)>1$, where $\lambda_{min}$ is the minimum non-zero value $\lambda$ could take.
As discussed above, heterogeneity of priors introduces the trade-off between persuading skeptics and dissuading believers, and one might expect that use of network signals and possibility of seeding could alleviate this trade-off. For instance, if a signal persuades both types of receivers, ideally the sender wants believers not to observe it, and skeptics to observe it very often. Is there a way of choosing seeds to achieve this?
By Lemma 2, for a signal to reach a large fraction of the population, we need the signals to be viral, i.e. the seeds of the signal needs to be on the relevant giant component. However, once the signals get viral, it is difficult to control who the signal reaches in a way that benefits the sender, unless the network exhibits certain features explored in the next subsection. In particular, in the baseline network, Lemma 5 states that once a signal gets viral, it must reach believers and sceptics with the same probability. Therefore, network signals cannot make sceptics observe informative signals more often than believers.
The reason for why sender could prefer network signals for some concave payoff function is as explained in the introduction. With public signals, all receivers of the same type must be choosing the same action. With network signals, receivers of the same type can have different observations and choose different actions. Roughly speaking, public signals will persuade either all receivers or no receivers (of the same type), whereas network signals allow the sender to always persuade some receivers. A sender who is very risk averse prefers the latter. The importance of the sender’s risk attitude has been noted by Arieli and Babichenko (2019). They relate optimality of public signals to the core of cooperative games. Their results imply that if receivers have the same prior and sender’s payoff is convex, then sender’s optimal payoff can be achieved using public signals. In contrast with this paper, they also find that when the sender’s payoff is strictly increasing [^11] Their condition is actually weaker. They only need the sender to prefer 1 receiver choosing action 1 to no receiver choosing action 1, public signals cannot be optimal if receivers have heterogeneous priors, while the sender prefers public signals even with heterogeneous priors in Proposition 1. The difference is mainly caused by Arieli and Babichenko (2019) having no communication between receivers. Communication between receivers impose restrictions on correlations of observations between different receivers, which reduces the sender’s payoff using personalised signals.
I provide a rough sketch of the proof for point 1. Suppose for simplicity that sender’s signal structure $\pi(\cdot)$ uses only 2 non-empty signals $s$ and $s^{\prime}$, where $s$ persuades both types of receivers to take action 1 and $s^{\prime}$ persuades only the believers, i.e. $\frac{\pi(s|\omega=1)}{\pi(s|\omega=0)}>\frac{\mu_{l}(\omega=0)}{\mu_{l}(\omega=1)}$, and $\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}<\frac{\pi(s^{\prime}|\omega=1)}{\pi(s^{\prime}|\omega=0)}<\frac{\mu_{l}(\omega=0)}{\mu_{l}(\omega=1)}$.
By Lemma 2, if no seeds are on the relevant giant component, then very few people will observe it. Therefore, for the sketch here, suppose that the seeds for $s$ are always on $L_{1}$ and the seeds for $s^{\prime}$ are always on $\hat{L}_{1}$. Let $a^{*}(\emptyset;t,d,\Pi)$ denote receiver’s optimal action upon observing $\emptyset$, given receiver’s degree $d$, type $t$ and sender’s strategy $\Pi$.
When $n$ gets large, roughly a proportion $\gamma_{t}p_{d}^{t}$ of the population has type $t$ and degree $d$. Among them, roughly a proportion $\zeta(t,d)$ are on the giant component $L_{1}$. Therefore, when $s$ is realised, roughly $\zeta(t,d)$ of those with type $t$ and degree $d$ observe $s$, and others choose $a^{*}(\emptyset;t,d,\Pi)$. Also, roughly a proportion $\hat{\zeta}(t,d)$ of them have a neighbour on the giant component in the subnetwork of type $h$ nodes. Therefore, when $s^{\prime}$ is realised, roughly $\hat{\zeta}(t,d)$ of those with type $t$ and degree $d$ observe $s^{\prime}$, and others choose $a^{*}(\emptyset;t,d,\Pi)$. By Lemma 5, $\zeta(t,d)$ and $\hat{\zeta}(t,d)$ don’t depend on $t$, so I omit the argument $t$ in this sketch. Also, $p_{d}^{h}$ and $p_{d}^{l}$ are the same in the baseline network so I omit the superscript. Using these, sender’s payoff is close to
$\sum_{\omega}\mu_{s}(\omega)\Bigg{[}\pi(s|\omega)v(\sum_{d}p_{d}\sum_{t}\gamma_{t}\big{[}\zeta(d)+(1-\zeta(d))a^{*}(\emptyset;t,d,\Pi)\big{]})\\ +\pi(s^{\prime}|\omega)v(\sum_{d}p_{d}\big{[}\gamma_{h}\hat{\zeta}(d)+(1-\hat{\zeta}(d))\sum_{t}\gamma_{t}a^{*}(\emptyset;t,d,\Pi)\big{]})+\pi(\emptyset|\omega)v(\sum_{d}p_{d}\sum_{t}\gamma_{t}a^{*}(\emptyset;t,d,\Pi))\Bigg{]}$ (1)
If $v$ is convex, then this is smaller than $\sum_{d}p_{d}V(d,\Pi)$ defined as
$\sum_{d}p_{d}\sum_{\omega}\mu_{s}(\omega)\Bigg{[}\pi(s|\omega)\zeta(d)v(1)+\pi(s^{\prime}|\omega)\hat{\zeta}(d)v(\gamma_{h})\\ +(1-\pi(s|\omega)\zeta(d)-\pi(s^{\prime}|\omega)\hat{\zeta}(d))v(\sum_{t}\gamma_{t}a^{*}(\emptyset;t,d,\Pi))\Bigg{]}\equiv\sum_{d}p_{d}V(d,\Pi)$ (2)
Therefore, to show that public signals are better, it’s sufficient to show that we can imitate the payoffs $V(d,\Pi)$ for each $d$. Consider a public signal strategy that sends the signal $s$ with probability $\pi(s|\omega)\zeta(d)$ and $s^{\prime}$ with probability $\pi(s^{\prime}|\omega)\hat{\zeta}(d)$ in each state $\omega$, and sends $\emptyset$ with the remaining probability. Note that these are roughly the probabilities that agents with degree $d$ observe $s$, $s^{\prime}$ and $\emptyset$ under the network signal strategy $\Pi$, therefore posteriors for each observation are also close to those under $\Pi$. A conjecture is that this implies that receiver’s action given each observation will be the same as those with degree $d$ under network signal strategy $\Pi$. A technical difficulty is that receiver’s action changes discontinuously at posterior 0.5, so posterior being close does not directly imply that they choose the same action. This is dealt with in the proof, and here for the sake of illustration let us take this conjecture to be true. Given this conjecture, one can check that sender’s payoff from this public signal structure is $V(d,\Pi)$.
Intuitively, given any $d$, the sender can use public signals to imitate the probabilities that receivers observe each signal under some network signal structure. This is possible because, conditional on degree, probabilities of observing signals are (roughly) the same for both types of receivers.
4.2 Homophily and type-dependent degree distribution
In this subsection, I study networks that could have homophily and correlation between degrees and types. The conditions for the results in this subsection will be stated in terms of $\{p_{d}^{t}\}_{d}$ and $q_{t}$, i.e. the degree distribution and expected proportion of neighbours of the same type, instead of the distribution over $\lambda$ and $q$. There are two reasons for this. Firstly, from data, it is usually easy to calculate the degree distribution and $q_{t}$, but it is difficult if not impossible to estimate $q$ and the distribution of $\lambda$. Therefore, stating the results in terms of the degree distributions and $q_{t}$ makes it easier to check whether the conditions of the propositions are satisfied. Secondly, all the results would go through when using the more general random networks model from Sadler (2020), which directly imposes arbitrary degree distributions and values of $q_{t}$, so the results are still relevant even we face some degree distribution that could not be generated by the more restrictive model used here. I will also discuss relations between degree distribution, $q_{t}$ and distribution over $\lambda$ and $q$.
I need to introduce some definition to state the results. First, I define forward (degree) distribution. The forward distribution captures the degree distribution of a randomly selected neighbour.
Definition 1.
The forward distribution $\{\tilde{p}_{d}^{t}\}_{d}$ for type $t$ is defined as $\tilde{p}_{d}^{t}\equiv\frac{p_{d+1}^{t}(d+1)}{\sum_{d}dp_{d}^{t}}$
Next I define what is meant when I say one type is more connected than the other
Definition 2.
I say that type $t$ receivers are weakly more connected than type $t^{\prime}$ if one of the two following conditions hold
  1. 1.
    The degree distributions $\{p_{d}^{t}\}_{d}$ and $\{p_{d}^{t^{\prime}}\}_{d}$ are the same.
  2. 2.
    $\{p_{d}^{t}\}_{d}$ FOSD (first order stochastically dominates) $\{p_{d}^{t^{\prime}}\}_{d}$. If $q_{t}\neq 1-q_{t^{\prime}}$, then we also require forward distribution of type $t$ to FOSD forward distribution of type $t^{\prime}$.
Like in the last subsection, we first study the probability that receivers observe viral signals.
Lemma 6.
If $q_{h}$ is (strictly) larger than $1-q_{l}$, then $\hat{\zeta}(h,d)$ is (strictly) larger than $\hat{\zeta}(l,d)$ for all $d$. If type $h$ receivers are weakly more connected than type $l$ receivers, then $\zeta(h,d)\geq\zeta(l,d)$ for all $d$.
Before stating the main proposition, I briefly comment on how the measure $q_{t}$ are related to measures of homophily. Coleman (1958)’s measure of homophily in this model is $\frac{q_{t}-\gamma_{t}}{1-\gamma_{t}}$, and there’s homophily for type $t$ if this measure is positive. Therefore, the network exhibits homophily if $q_{t}>\gamma_{t}$ for both types. Note that if $q_{h}>\gamma_{h}$ and $q_{l}>\gamma_{l}$, then we must have $q_{h}>1-q_{l}$ since $\gamma_{h}$ and $\gamma_{l}$ sum up to 1. Therefore, the condition in the lemma 6 and the proposition below are weaker than requiring the network to exhibit homophily.
Using lemma 6, we can show that conclusions from proposition 1 extend even when the network exhibits homophily, if the believers are weakly more connected than the sceptics.
Proposition 2.
If the type $h$ receivers are weakly more connected than type $l$ receivers and $q_{h}\geq 1-q_{l}$, conclusions of proposition 1 still hold, i.e.
  1. 1.
    If $v(.)$ is linear or convex, $\limsup_{n\rightarrow\infty}V^{*}_{net}(n)\leq V^{*}_{pub}$.
  2. 2.
    If $\hat{c}>0$, there exists some strictly concave $v(.)$ such that $\liminf_{n\rightarrow\infty}V^{*}_{net}>V^{*}_{pub}$.
Remark 2 also holds so is not stated again for the sake of space.
Homophily is often cited as one of the reasons why advertising and political campaigns on social media is more effective than traditional forms of advertisement, since it influences what information people could observe. Lemma 6 indeed shows that homophily and type-degree correlation introduce asymmetries between the two groups. Roughly speaking, the sender would like to send signals more often to sceptics and keep information away from believers. From Lemma 6, when believers are weakly more connected in the network, the asymmetry introduced is the opposite of what the sender desires.
To discuss the mechanism in more detail, I again impose the simplifying assumption that the sender uses only two signals $s$ and $s^{\prime}$, with $s$ persuading both types and $s^{\prime}$ persuading believers only. From Lemma 6 and focusing on the case where the two types are equally connected, homophily makes believers observe signal $s^{\prime}$ more often than sceptics. This asymmetry means that the imitation argument from last subsection does not directly work. Nonetheless, we could still imitate the probabilities of observing each signals for receivers of a given type and degree $d$. For any degree $d$, there are two cases. If sceptics with degree $d$ choose action 0 upon observing $\emptyset$, then they only choose action 1 when they see $s$. Therefore, if we let them observe signals with the same probabilities as believers with degree $d$, then their probabilities of choosing action 1 will not change. In the second case, suppose sceptics choose action 1 upon observing $\emptyset$, then believers must always choose action 1 when they observe signals with the same probabilities as sceptics since they have higher priors. Therefore, when the sender has convex payoff functions, imitating probabilities for one of the types gives the sender higher payoff than her ’payoff’ from type $d$ receivers (i.e. counterpart here for $V(d,\Pi)$ in last subsection).
Now, if believers are in addition more connected, then they will also observe $s$ with higher probabilities. This is exactly the opposite of what the sender desires. The sender would like sceptics to observe $s$ since they need convincing, and keep $s$ away from believers since observing $s$ too often can make $\emptyset$ too informative of the bad state.
Note that this result does not say that homophily is not useful for the sender. Homophily makes believers more clustered together, and signals that only persuade believers can spread more widely. Compared to a network without homophily, sender could send these signals less often while still maintaining the same probability that a believer with degree $d$ observes them. These signals are realised less often now, but are observed by more receivers when realised, which increases risk and potentially makes sender better off.
The next remark relates the degree distributions and $q_{t}$ to the distributions of $\lambda$ and $q$.
Remark 3.
  1. 1.
    $q_{h}>1-q_{l}$ if and only if $q<1$, and $q_{h}=1-q_{l}$ if and only if $q=1$.
  2. 2.
    If $q=1$, type $t$ is weakly more connected than type $t^{\prime}$ if $f_{t}(\cdot)$ FOSD $f_{t^{\prime}}(\cdot)$ or $f_{t}$ does not depend on $t$.
  3. 3.
    Given any $f_{t^{\prime}}(\cdot)$, we can find some $r$ large enough such that type $t$ is more connected than type $t^{\prime}$ if $f_{t}(r\lambda)=f_{t^{\prime}}(\lambda)$ for all $\lambda$.
    A special case is the island model often used in literature where all receivers have the same $\lambda$. In these models, the larger group is more connected.
Finally, I present conditions under which information can be sent to sceptics more often than believers, and therefore network signals are better than public signals even for risk loving sender.
Proposition 3.
Suppose that there’s no type $l$ receiver with $\lambda$ smaller than some arbitrary $\lambda^{\prime}>0$. We can find some $\overline{x}$ large enough such that if expected degree of type $l$ receivers is larger than $\overline{x}$ and expected degree of type $h$ receivers is smaller than $1/\overline{x}$, then $\liminf_{n\rightarrow\infty}V_{net}^{*}(n)>V_{pub}^{*}$.
Intuitively, believers are less connected than sceptics, and therefore they receive information less often than sceptics, which is desirable for the sender. As an extreme example, consider a network where all believers are isolated and all sceptics are connected together. This allows the sender to achieve the ideal outcome of sending information only to sceptics and keeping believers completely uninformed.
The high level intuition underlying this and the last subsection is that, it is difficult to control who sees viral signals using seeding. It is the network structures and how people share signals that determine who observe viral signals. Therefore, whether the sender prefers network signals depends on, roughly speaking, whether the network structure causes believers to receive signals less often than sceptics. From the discussion following proposition 2, it is mainly the probability of observing signals that persuade both types that matters. This is in turn influenced by type-degree correlation. Therefore, whether the sender prefers network signals is mainly determined by whether sceptics are more connected, and homophily cannot reverse the sender’s preference for public signals over network signals when believers are more connected.
4.3 Voting game
An often studied persuasion game is the voting game. In voting games, the sender only cares about whether the proportion of receivers taking sender’s preferred action reaches a certain threshold. Let the sender’s payoff function be defined in the following way: There exists some $\overline{x}<1$ such that $v(x)=1$ if $x\geq\overline{x}$ and $v(x)=0$ otherwise. To make the problem non-trivial, I assume that $\overline{x}>\gamma_{h}$.
To state the result, I introduce the following definition. Define $d_{vote}$ to be the smallest $d$ such that
$\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}\hat{\zeta}(l,d)}{1-\hat{\zeta}(l,d)}\geq 1$
Let $d_{vote}$ be $\infty$ if there exists no such $d$. As shown in proof of the next proposition, if a type $l$ receiver has degree lower than $d_{vote}$, then she cannot choose action 1 upon observing $\emptyset$. Using this definition, we can provide a necessary condition and a sufficient condition for network signals to outperform public signals.
Proposition 4.
A necessary condition for $\limsup_{n\rightarrow\infty}V_{net}^{*}(n)>V_{pub}^{*}$ is
$\sum_{d\geq d_{vote}}\gamma_{l}p_{d}^{l}\geq\overline{x}-\gamma_{h}$
A sufficient condition for $\liminf_{n\rightarrow\infty}V_{net}^{*}(n)>V_{pub}^{*}$ is
$\sum_{d\geq d_{vote}+1}\gamma_{l}p_{d}^{l}(1-\hat{\zeta}(l,d))>\overline{x}-\gamma_{h}$
Roughly speaking, both conditions say that network signals perform better when $\overline{x}-\gamma_{h}$ is small relative to the proportion of type $l$ receivers who could choose action 1 upon observing $\emptyset$. Sender’s payoff function $v$ in voting games is neither concave nor convex, so results in previous sections do not directly apply, but they are still useful to get some intuition for proposition 4. Consider a reduced form problem where sender chooses between always getting moderate levels of support from skeptics, or sometimes getting a lot of support from skeptics and sometimes losing a lot of support from believers. The former option involves less risk. When $\gamma_{h}$ is small and threshold $\overline{x}$ is large, the sender prefers the latter option, since getting moderate support is not enough to reach the threshold. The sender behaves as if she is risk loving, and results in previous subsections suggest that in these cases public signals are better. When $\gamma_{h}$ is close to $\overline{x}$, sender prefers the former option since getting way above the threshold is not valuable. The sender behaves as if she is risk averse, and previous results suggest that network signals can be better.
If skeptics do not choose action 1 upon observing $\emptyset$, they choose action 1 only upon observing signals that persuade both types of receivers. The sender can only get enough support if a large enough fraction of sceptics observe such signals. Therefore, the sender would like to maximise the probability of sending such signals and maximise the proportion of people they reach. This is achieved by using public signals. This explains the importance of having skeptics choosing action 1 upon seeing $\emptyset$ for network signals to be preferred to public signals.
5 Extensions and Discussion
5.1 Alternative sharing rules
As discussed in detail above, the key mechanism behind all the results on public signals being better is that believers observe signals more often than sceptics. These results would still go through with any other rule where (conditional on degree) believers share each signal $s$ more often than sceptics. Such sharing rules ensure that, under conditions stated in the proposition 1 and 2, believers observe each signal more often than sceptics. Note that the sharing rule studied above falls into this more general class, so the proofs are very similar.
I now discuss one sharing rule that does not fall into the class above, but the results in the paper still go through. Suppose that now receivers share a signal $s$ if their action ’agrees’ with content of the signal. Formally, they share the signal if their action is 1 and the signal is more likely in state 1, or if their action is 0 and the signal is more likely in state 0. All receivers will pass on signals that persuade both types and neither types. Only sceptics pass on signals that persuade only the believers, so sceptics can observe these signals more often than believers even if believers are weakly more connected.
However, this cannot make network signals better than public signals if the believers are more connected. To see this, one can go back to the discussion after proposition 2, explaining why public signals are still preferable when network exhibits homophily. Note that the argument there does not use the fact that signals persuading only believers are observed more often by believers. More formally, Lemma 8 in the appendix which presents the formal version of imitation arguments in the main text, does not use that signals persuading only believers are observed more often by believers. Therefore, all results would go through in essentially the same way with this alternative sharing rule.
5.2 Non-vanishing fraction of seeds and optimal network signals
The assumption of seeds being a vanishing fraction of population can be relaxed as long as the seeds only take a fraction of population that is small enough. An upper bound on sender’s payoff can be obtained by assuming that all the seeds and others on the same component with them, excluding those on the giant component, always choose action 1 and remove them from the network. We can then analyze the remaining network and results above will go through. This implies that results would go through as long as the fraction of seeds is small enough even though it is non-vanishing.
On the other hand, if the fraction of seeds is larger than $\gamma_{l}$, i.e. the fraction of sceptics. Then the sender could always just directly send signals that only realise in state 0 to all the sceptics, to persuade them to take action 1 upon observing $\emptyset$. Believers will receive no information. Sender achieves her highest possible payoff (when $v$ is linear or convex) by doing this.
For intermediate sizes of seed set, the analysis is more complicated. The main reason is that we now need to consider components other than the giant component. These components are less well studied. Actually, there exist no closed form solutions even for probabilities of being on giant component (i.e. the functions $\zeta$ and $\hat{\zeta}$). Therefore, it is actually difficult to even explicitly determine what receivers do upon observing $\emptyset$ and thus sender’s payoff from a given strategy. This is why, as illustrated above, most results in this paper are obtained from bounding payoffs of network signals from above or below, usually by some imitation argument. Although it would be very interesting to look at sender’s optimal network signal, it is very challenging, since it is not easy to even determine payoff for a given strategy. Introducing intermediate sizes of seed set will make these difficulties more pronounced.
5.3 More than one signal spreading on network
The assumption of having 1 signal spreading on the network is mainly made for tractability. Without this assumption, there may not even exist an equilibrium given some strategies of the sender. To see this, consider an example where there are 2 receivers connected to each other. One of them is a sceptic and the other one is a believer. Suppose that the sender sends a signal that persuades both groups (if it’s the only signal observed) to the sceptic, and a signal that persuades neither to the believer. Also suppose that observing both persuades the believer but not the sceptic. In this case, if sceptic chooses action 1 and shares the signal to the believer, then the believer must also choose action 1. However, this passes on the bad news to the sceptic, which means the sceptic should actually take action 0, and her good news will not be shared in the first place.
One reason this issue arises is that the current model does not explicitly consider time, and assumes that sharing information is simultaneous and instantaneous. This is not an issue with just one signal. With only 1 signal, upon observing some signal, receivers know that they cannot obtain any further information and their posterior will not change. Therefore, we lose no insights by abstracting away from time. However, as illustrated above, we need to introduce time when there are multiple signals. This further introduces the complication of how receivers share information, anticipating that their beliefs may change in the future. The current model is silent on this. This is an interesting direction for future research.
6 Conclusion
There has been a lot of discussion on exploiting network information for advertising and political campaigns. This paper studies whether and when using network signals outperforms public signals. In this paper, receivers have heterogeneous priors and pass on a signal to their neighbours if they are persuaded by the signal. This paper finds that risk netural and risk loving senders prefer public signals, if believers are more connected than sceptics. This holds even if the network exhibits homophily. The sender may prefer network signals if she is risk averse or scceptics are sufficiently more connected than believers.
The main intuition underlying results in this paper is that, the sender needs the signal to get viral in order to have an impact. However, it is difficult to control who can observe viral signals using seeding. The sender would like to send more information to sceptics than believers, but this is only possible when sceptics are more connected than believers.
In many situations, the campaigner has access to tools that can send public signals. The results in this paper imply that, in many cases, resources may be more efficiently spent on sending public signals, rather than on trying to understand network structures and picking the optimal seeds. In addition, persuasion on large networks is complicated, while public signals have been well studied. From these comparisons, we can use sender’s value from public signals to obtain what sender could get at most or at least using network signals. Even if the sender has no access to public signals, the results could be useful to judge whether a campaign using network signals will be profitable for the sender.
Appendix A Appendix: Proofs
I use $PO(d,x)$ to denote the probability that a Poisson random variable with intensity $x$ is equal to $d$. I define $D(t,\lambda)\equiv\lambda(\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda))$ where $t^{\prime}\neq t$. This captures the expected degree of node with type $t$ and connectedness $\lambda$. The next subsection proves the main results, and the subsection after that proves the technical lemmas in Section 3.
A.1 Proof of main results
Using the lemmas 2, 3 and 4, I can now approximate probabilities that receivers observe an non-empty signal $s\neq\emptyset$. Let $C(.)$ and $\hat{C}(.)$ be functions over the set of seeds and networks defined in the following way
$C(z,g^{n})=\begin{cases}1&\text{ if }z\cup L_{1}\neq\emptyset\\ 0&\text{otherwise}\end{cases}\qquad\hat{C}(z,g^{n})=\begin{cases}1&\text{ if }z\cup\hat{L}_{1}\neq\emptyset\\ 0&\text{otherwise}\end{cases}$
In words, $C(z,g^{n})$ is 1 if seeds are on giant component, and $\hat{C}(z,g^{n})$ is 1 if seeds are on giant component after deleting type $l$ nodes. Define $\mathbf{A}^{n}(\epsilon,\overline{d})$ to be the set of networks with $n$ receivers that satisfy the following properties: Given any sender’s strategy $\Pi^{n}$ in the network signal game with $n$ receivers, we have
  1. 1.
    Let $s\neq\emptyset$ be a signal that persuades both types of receivers according to $\Pi^{n}$. Conditional on $g^{n}\in\mathbf{A}^{n}(\epsilon,\overline{d})$, signal $s$ and seeds $z$ being realised, the proportion of receivers with type $t$ and degree $d\leq\overline{d}$ who observe $s$ is at most $\epsilon$ from $C(z,g^{n})\zeta(t,d)$.
  2. 2.
    Let $s^{\prime}$ be a signal that persuades only type $h$ receivers according to $\Pi^{n}$. Conditional on $g^{n}\in\mathbf{A}^{n}(\epsilon,\overline{d})$, signal $s^{\prime}$ and seeds $z$ being realised, the proportion of receivers with type $t$ and degree $d\leq\overline{d}$ who observe $s^{\prime}$ is at most $\epsilon$ from $\hat{C}(z,g^{n})\hat{\zeta}(t,d)$.
  3. 3.
    If $s^{\prime\prime}$ is a signal that persuades no receiver according to $\Pi^{n}$, the proportion of receivers who observe $s^{\prime\prime}$ is at most $\epsilon$.
Lemma 7.
Given any $\overline{d}\in\mathbb{N}$ and $\epsilon>0$, there exists an $n_{1}$ large enough such that for any $n\geq n_{1}$, $Pr(g^{n}\in\mathbf{A}_{n}(\epsilon,\overline{d}))>1-\epsilon$
Intuitively, this lemma states that as $n$ gets large, the probability that a receiver observes a certain signal is given by the probability that it hits and the receiver is on the relevant giant component.
Proof.
By lemmas 3 and 4, for any $\epsilon>0$, we can find $n$ large enough such that with arbitrarily high probability, the proportion of receivers with type $t$ and degree $d$ who observe $s$ conditional on seeds being on relevant giant component is within $[\zeta(t,d)+2\epsilon,\zeta(t,d)-2\epsilon]$ if $s$ persuades both types of receivers, and is within $[\hat{\zeta}(t,d)+2\epsilon,\hat{\zeta}(t,d)-2\epsilon]$ if $s$ persuades only type $h$ receivers. Result then follows from definition of functions $C(.)$ and $\hat{C}(.)$, and lemma 2. ∎
Lemma 5 and proposition 1 are special cases of lemma 6 and proposition 2, so I only prove the latter two.
Proof of lemma 6:
The first part of the result directly follows from the proof of lemma 14.
For the second part, by definition, we have that $\rho(\infty|t,\lambda)=\sum_{d}Poisson(d;D(t,\lambda))\rho(\infty|t,d)$ From remark 4, we have
$1-\rho(\infty|t,d)=(q_{t}\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{t}(\lambda^{\prime})(1-\rho(\infty|t,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{t}(\lambda^{\prime})}+(1-q_{t})\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{t^{\prime}}(\lambda^{\prime})(1-\rho(\infty|t^{\prime},\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{t^{\prime}}(\lambda^{\prime})})^{d}$ (A.1)
This directly implies the result when $q_{h}=1-q_{l}$. I now discuss the case with $q_{h}\neq 1-q_{l}$. The forward distribution for type $t$ is given by
$\frac{(d+1)\sum_{\lambda}f_{t}(\lambda)Poisson(d+1;\lambda(\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda)))}{\sum_{\lambda}\lambda f_{t}(\lambda)(\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda))}$
which by the probability mass function of Poisson distribution is equal to
$\frac{\sum_{\lambda}\lambda f_{t}(\lambda)Poisson(d;\lambda(\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda)))}{\sum_{\lambda}\lambda f_{t}(\lambda)}=\frac{\sum_{\lambda}\lambda f_{t}(\lambda)Poisson(d;D(t,\lambda))}{\sum_{\lambda}\lambda f_{t}(\lambda)}$
We can substitute $\rho(\infty|t,\lambda)=\sum_{d}Poisson(d;D(t,\lambda))\rho(\infty|t,d)$ into RHS of Equation A.1, which along with the expression for forward distribution gives
$1-\rho(\infty|t,d)=(q_{t}(1-\sum_{d}\tilde{p}_{d}^{t}\rho(\infty|t,d))+(1-q_{t})(1-\sum_{d}\tilde{p}_{d}^{t^{\prime}}\rho(\infty|t^{\prime},d)))^{d}$
Suppose that $1-\rho(\infty|h,d)>1-\rho(\infty|l,d)$ for some $d$, then from the expression above, it must also be the case for all $d$. In particular, it must be true for $d=1$. Also, let $\frac{1-\rho(\infty|h,1)}{1-\rho(\infty|l,1)}=\alpha$. Notice that $1-\rho(\infty|t,d)=(1-\rho(\infty|t,1))^{d}$, and we have $\frac{1-\rho(\infty|h,d)}{1-\rho(\infty|l,d)}=\alpha^{d}$. Therefore, we can write
$\alpha(1-\rho(\infty|l,1))=q_{h}(\sum_{d}\tilde{p}_{d}^{h}\alpha^{d}(1-\rho(\infty|l,1))^{d})+(1-q_{h})\sum_{d}\tilde{p}_{d}^{l}(1-\rho(\infty|l,1))^{d}$
$1-\rho(\infty|l,1)=q_{l}(\sum_{d}\tilde{p}_{d}^{l}(1-\rho(\infty|l,1))^{d})+(1-q_{l})(\sum_{d}\tilde{p}_{d}^{h}\alpha^{d}(1-\rho(\infty|l,1))^{d})$
I now show that there can exist no $\rho(\infty|l,1)$ and $\alpha>1$ where the two equalities hold at the same time. I separate this into two case. If $q_{h}<1-q_{l}$, then $\alpha>1$ implies that RHS of the first equation is smaller than that of the second, which contradicts with $\alpha>1$.
If $q_{h}>1-q_{l}$, for $\alpha$ to be greater than 1, we must have $\sum_{d}\tilde{p}_{d}^{h}\alpha^{d}(1-\rho(\infty|l,1))^{d}>\sum_{d}\tilde{p}_{d}^{l}(1-\rho(\infty|l,1))^{d}$. Therefore, we must choose some $1-\rho(\infty|l,1)>\sum_{d}\tilde{p}_{d}^{l}(1-\rho(\infty|l,1))^{d}$, otherwise the second equation can’t hold.
Given such $1-\rho(\infty|l,1)$, if $\alpha=1$, the LHS of both equations must be larger than the RHS. The largest value of $\alpha$ is when $\alpha(1-\rho(\infty|l,1))=1$. WHen $\alpha$ takes this value, the LHS of the first equation is 1 and RHS is less than 1. Also, notice that the LHS of the first equation is linear in $\alpha$ while the RHS is convex. These facts imply that LHS of the first equation must always be larger than the RHS for all $\alpha\geq 1$. This concludes the proof. ∎
For any network signal strategy $\Pi^{n}=(\pi,Z)$ in the game with $n$ receivers, let $a^{*}(s,t,d;\Pi^{n})$ denote optimal action of receivers with type $t$ and degree $d$ when they observe $s$, for $s\in S\cup\emptyset$. Let $N_{t,d,s}(g^{n},z)$ denote the number of receivers of type $t$ and degree $d$ who observe signal $s$ given network $g^{n}$ and seeds $z$. Define function $\hat{V}(d,\Pi^{n})$ to be
$\hat{V}(d,\Pi^{n})\equiv\sum_{\omega}\mu_{s}(\omega)\sum_{s\in S\cup\{\emptyset\}}\pi(s|\omega)\bigg{[}\sum_{g^{n}}Pr(g^{n})\\ v\big{(}\sum_{t}\gamma_{t}\frac{N_{t,d,s}(g^{n},z)}{N_{t,d}}a^{*}(s,t,d;\Pi^{n})+\sum_{t}\gamma_{t}(1-\frac{N_{t,d,s}(g^{n},z)}{N_{t,d}})a^{*}(\emptyset,t,d;\Pi^{n})\big{)}\bigg{]}$ (A.2)
Using this function, the next lemma formally presents (part of) the imitation argument illustrated in main text.
Lemma 8.
Suppose that $\zeta(d,h)\geq\zeta(d,l)$. If $v$ is convex or linear, for any $\epsilon>0$ and $d$, there exists some $n_{1}$ large enough such that for all $n\geq n_{1}$, we have $\hat{V}(d,\Pi^{n})\leq V_{pub}^{*}+\epsilon$ for all possible network signal strategy $\Pi^{n}$.
Proof.
For any $\beta^{\prime}>0$, we can find a $\beta>0$ such that $sup(v(x+2\beta)-v(x))<\beta^{\prime}$. For this $\beta$, by lemma 7 we can find some $n_{1}$ such that for all $n>n_{1}$, we have $Pr(g^{n}\in\mathbf{A}_{n}(\beta,\overline{d}))>1-\beta$. Let $S_{good}$ denote set of signals in $S$ that persuade both types of receivers, $S_{int}$ denote set of signals in $S$ that persuade only type $h$ receivers and $S_{bad}$ denote those that persuade neither group. For $j\in\{good,int,bad\}$, let
$\hat{\pi}_{j}(t,\omega,\Pi^{n})\equiv\sum_{s\in S_{j}}\pi(s|\omega)\sum_{g^{n}}Pr(g^{n})\frac{N_{t,d,s}(g^{n},z)}{N_{t,d}}$
We consider 2 cases. First, suppose that $a^{*}(\emptyset,l,d;\Pi^{n})=0$. We know that $a^{*}(s,l,d;\Pi^{n})\geq a^{*}(\emptyset,l,d;\Pi^{n})$ for all $s\in S$. By definition of $\mathbf{A}_{n}(\beta,\overline{d})$, we know that $\frac{N_{l,d,s}(g^{n},z)}{N_{l,d}}-\frac{N_{h,d,s}(g^{n},z)}{N_{h,d}}$ is smaller than $2\beta$ if $g^{n}\in\mathbf{A}_{n}(\beta,\overline{d})$ and $s\in S_{good}$. Using these facts, how $n_{1}$, $\beta$ are chosen and convexity of $v$, $\hat{V}(d,\Pi^{n})$ is smaller than
$\sum_{\omega}\mu_{s}(\omega)\bigg{[}\bigg{(}1-\sum_{j\in\{good,int,bad\}}\hat{\pi}_{j}(h,\omega,\Pi^{n})\bigg{)}v\big{(}\gamma_{h}a^{*}(\emptyset,t,d;\Pi^{n})\big{)}\\ +\hat{\pi}_{good}(h,\omega,\Pi^{n})v(1)+\hat{\pi}_{int}(h,\omega,\Pi^{n})v(\gamma_{h})\bigg{]}+v(1)\beta+\beta^{\prime}$ (A.3)
Consider sender’s public signal strategy $\pi_{2}$ described as follows. For $j\in\{good,int,bad\}$
$\pi_{2}(s_{j}|\omega)=\hat{\pi}_{j}(h,\omega,\Pi^{n})\qquad\pi_{2}(\emptyset|\omega)=1-\sum_{j\in\{good,int,bad\}}\hat{\pi}_{j}(h,\omega,\Pi^{n})$
By lemma 1 and construction of $\pi_{2}$, all receivers choose action 1 upon observing signal $s_{good}$ with the public signal structure $\pi_{2}$. Similarly, type $h$ receivers choose action 1 upon observing $s_{int}$ and no receiver choose action 1 upon observing $s_{bad}$.
Also, receivers observe $\emptyset$ in state $\omega$ with the same probability as receivers with type $h$ and degree $d$ under network signal structure $\Pi^{n}$. Therefore, type $h$ receivers choose $\alpha(\emptyset,h,d;\Pi^{n})$ upon observing $\emptyset$. Therefore, sender’s payoff using $\pi_{2}$ (as $n$ goes to infinity) is expression A.3 minus $\beta v(1)+\beta^{\prime}$. Since sender’s limiting payoff from $\pi_{2}$ is smaller than $V_{pub}^{*}$ and $\beta,\beta^{\prime}$ can be arbitrarily small, the result follows for this case.
In second case, suppose $\alpha(\emptyset,l,d;\Pi^{n})=1$. We know that $\alpha(s,h,d;\Pi^{n})\leq 1$ for all $s$. Using this fact and convexity of $v$, expression $\hat{V}(d,\Pi^{n})$ is smaller than
$\beta v(1)+\beta^{\prime}+\sum_{\omega}\mu_{s}(\omega)\bigg{[}\bigg{(}1-\sum_{j\in\{good,int,bad\}}\hat{\pi}_{j}(l,\omega,\Pi^{n})\bigg{)}v\big{(}1\big{)}\\ +\hat{\pi}_{good}(l,\omega,\Pi^{n})v(1)+\hat{\pi}_{int}(l,\omega,\Pi^{n})v(\gamma_{h})\bigg{]}\equiv\hat{V}^{l}(d,\Pi^{n})+\beta v(1)+\beta^{\prime}$ (A.4)
Consider sender’s public signal strategy $\pi_{3}$ described as follows. For $j\in\{good,int,bad\}$
$\pi_{3}(s_{j}|\omega)=\hat{\pi}_{j}(l,\omega,\Pi^{n})\qquad\pi_{3}(\emptyset|\omega)=1-\sum_{j\in\{good,int,bad\}}\hat{\pi}_{j}(h,\omega,\Pi^{n})$
By essentially the same argument as in the previous case, all receivers choose action 1 upon observing signal $s_{good}$ with the public signal structure $\pi_{3}$. Similarly, type $h$ receivers choose action 1 upon observing $s_{int}$ and no receiver choose action 1 upon observing $s_{bad}$.
Also, receivers observe $\emptyset$ in state $\omega$ with the same probability that a receiver with type $l$ and degree $d$ observes $\emptyset$ under network signal structure $\Pi^{n}$. Therefore, type $l$ receivers choose action 1 upon observing $\emptyset$, and type $h$ receivers must also choose action 1 upon observing $\emptyset$. Sender’s payoff from using public signals $\pi_{3}$ (as $n$ goes to infinity) is $\hat{V}^{l}(d,\Pi^{n})$, which is expression A.4 minus $\beta v(1)+\beta^{\prime}$. Since sender’s limiting payoff from $\pi_{3}$ is smaller than $V_{pub}^{*}$, and $\beta,\beta^{\prime}$ can be arbitrarily small, the result follows.
The proof above immediately implies the following result.
Corollary 2.
If $\alpha(\emptyset,l,d;\Pi^{n})=1$, then $V_{pub}^{*}\geq\hat{V}^{l}(d,\Pi^{n})$, where $\hat{V}^{l}(d,\Pi^{n})$ is defined in expression A.4.
Proof of Proposition 2 point 1:
For any $\beta^{\prime}>0$, we can find a $\beta>0$ such that $sup(v(x+8\beta)-v(x))<\beta^{\prime}$. For this $\beta$, we can find some $\overline{d}$ such that $\sum_{d>\overline{d}}(p_{d}^{h}+p_{d}^{l})<\beta$. Fix this $\overline{d}$. By Lemma 7, Lemma 8 and Remark 1, we can find some $n_{1}$ such that for all $n>n_{1}$, we have $\hat{V}(d,\Pi^{n})\leq V_{pub}^{*}+\beta^{\prime}$ for all $\Pi^{n}$ and the following conditions are satisfied with probabilty at least $(1-\beta)$:
$\sum_{d>\overline{d}}\frac{N_{h,d}+N_{l,d}}{n}<\beta$, $\frac{N_{t,d}}{n}\in[\gamma_{t}p_{d}^{t}(1-\beta),\gamma_{t}p_{d}^{t}(1+\beta)]$ and $g^{n}\in\mathbf{A}_{n}(\beta,\overline{d})$ for all $d\leq\overline{d}$, $t\in\{h,l\}$.
Let $S_{good}$, $S_{int}$ and $S_{bad}$ have the same meaning as in proof of lemma 8. In the game with $n>n_{1}$ receivers, by condition 1, sender’s payoff from any $\Pi^{n}_{1}$ is smaller than,
$\beta v(1)+\sum_{\omega}\mu_{s}(\omega)\sum_{s\in S\cup\{\emptyset\}}\pi_{1}(s|\omega)\bigg{[}\sum_{g^{n}}Pr(g^{n})\\ v\big{(}2\beta+\sum_{t,d\leq\overline{d}}p_{d}^{t}\gamma_{t}\frac{N_{t,d,s}(g^{n},z)}{N_{t,d}}a^{*}(s,t,d;\Pi^{n}_{1})+p_{d}^{t}\gamma_{t}(1-\frac{N_{t,d,s}(g^{n},z)}{N_{t,d}})a^{*}(\emptyset,t,d;\Pi^{n}_{1})\big{)}\bigg{]}$ (A.5)
We consider two cases. In the first case, suppose that $a^{*}(\emptyset,l,d;\Pi_{1}^{n})=0$ for all $d\leq\overline{d}$. In this case, by definition, we also know that $a^{*}(s,l,d;\Pi_{1}^{n})=0$ for any signal $s\notin S_{good}$. Conditional on $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$, we know that $\frac{N_{l,d_{1},s}(g^{n},z)}{N_{l,d_{1}}}-\frac{N_{l,d_{2},s}(g^{n},z)}{N_{l,d_{2}}}\leq 2\beta$ for any $d_{1},d_{2}\leq\overline{d}$ and $d_{1}\leq d_{2}$. Since degree distribution for type $h$ FOSD that of type $l$, by how $\beta$ is chosen, condition 1 and convexity of $v$, A.5 is smaller than
$\sum_{d\leq\overline{d}}p_{d}^{h}\hat{V}^{n}(d,\Pi^{n}_{1})+\beta v(1)+\beta^{\prime}\leq V_{pub}^{*}+\beta v(1)+2\beta^{\prime}$
The result for this case then follows since $\beta$ and $\beta^{\prime}$ can be arbitrarily small.
Now consider the second case that there exists some $d$ such that $a^{*}(\emptyset,l,d;\Pi^{n}_{1})=1$. Let $d_{min}$ denote the smallest $d\leq\overline{d}$ such that $a^{*}(\emptyset,l,d_{min};\Pi^{n}_{1})=1$ and $a^{*}(\emptyset,l,d;\Pi^{n}_{1})=0$ for all $d<d_{min}$. Note that this implies $a^{*}(\emptyset,l,d_{min};\Pi_{1}^{n})\geq a^{*}(\emptyset,t,d;\Pi_{1}^{n})$ for all $d$, $t$. Again, if $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$, we know that $\frac{N_{t,d_{1},s}(g^{n},z)}{N_{t,d_{1}}}-\frac{N_{t,d_{2},s}(g^{n},z)}{N_{t,d_{2}}}\leq 2\beta$ for any $d_{1},d_{2}\leq\overline{d}$ and $d_{1}\leq d_{2}$. From these, we know that, for any $s$ and $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$,
$\sum_{d<d_{min}}p_{d}^{l}\gamma_{l}\frac{N_{l,d,s}(g^{n},z)}{N_{l,d}}a^{*}(s,l,d;\Pi^{n}_{1})+p_{d}^{l}\gamma_{l}(1-\frac{N_{l,d,s}(g^{n},z)}{N_{l,d}})a^{*}(\emptyset,l,d;\Pi^{n}_{1})\leq\\ 2\beta+\sum_{d<d_{min}}p_{d}^{h}\gamma_{l}\frac{N_{l,d,s}(g^{n},z)}{N_{l,d}}a^{*}(s,l,d;\Pi^{n}_{1})+p_{d}^{h}\gamma_{l}(1-\frac{N_{l,d,s}(g^{n},z)}{N_{l,d}})a^{*}(\emptyset,l,d;\Pi^{n}_{1})+\\ (\sum_{d<d_{min}}p_{d}^{l}\gamma_{l}-\sum_{d<d_{min}}p_{d}^{h}\gamma_{l})(\frac{N_{l,d_{min},s}(g^{n},z)}{N_{l,d_{min}}}a^{*}(s,l,d_{min};\Pi^{n}_{1})+(1-\frac{N_{l,d_{min},s}(g^{n},z)}{N_{l,d_{min}}}))$ (A.6)
Also, for all $d\geq d_{min}$, $t$, $s$ and $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$, we have
$p_{d}^{t}\gamma_{t}\frac{N_{t,d,s}(g^{n},z)}{N_{t,d}}a^{*}(s,t,d;\Pi^{n}_{1})+p_{d}^{t}\gamma_{t}(1-\frac{N_{t,d,s}(g^{n},z)}{N_{t,d}})a^{*}(\emptyset,t,d;\Pi^{n}_{1})\leq\\ p_{d}^{t}\gamma_{t}\frac{N_{l,d_{min},s}(g^{n},z)}{N_{l,d_{min}}}a^{*}(s,t,d;\Pi^{n}_{1})+p_{d}^{t}\gamma_{t}(1-\frac{N_{l,d_{min},s}(g^{n},z)}{N_{l,d_{min}}})+2\gamma_{t}p_{d}^{t}\beta$ (A.7)
By how $\beta$ is chosen, convexity of $v$ and the previous two inequalities, expression A.5 is smaller than
$\beta v(1)+\beta^{\prime}+\sum_{d<d_{min}}p_{d}^{h}\hat{V}(d,\Pi_{1}^{n})+\sum_{d_{min}\leq d\leq\overline{d}}p_{d}^{h}\hat{V}^{l}(d_{min},\Pi_{1}^{n})$
Since $\hat{V}(d,\Pi^{n}_{1})\leq V_{pub}^{*}+\beta^{\prime}$ and $\hat{V}^{l}(d_{min},\Pi_{1}^{n})\leq V_{pub}^{*}$ (by 2), the first point follows since $\beta$ and $\beta^{\prime}$ can be arbitrarily small. ∎
Proof of Proposition 2 point 2:
Let the sender use the following network signalling strategy for all $n$
$\pi_{1}(s^{\prime}|\omega=1)=r(1-\kappa)\qquad\pi_{1}(s^{\prime}|\omega=0)=\frac{\mu_{h}(\omega=1)}{\mu_{h}(\omega=0)}r(1-\kappa)\qquad\pi(\emptyset|\omega)=1-\pi_{1}(s^{\prime}|\omega)$
where $r\geq 1$ and $\kappa=(\frac{\mu_{h}(\omega=1)}{\mu_{h}(\omega=0)}-1)/(\frac{\mu_{h}(\omega=1)}{\mu_{h}(\omega=0)}-\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)})$. Sender always chooses one seed on $\hat{L}_{1}$. It can be verified that type $h$ receivers must lways choose action 1. Now, consider the posterior of a type $l$ receiver with degree $d$ upon observing $\emptyset$. Using lemma 7, we know that it converges to:
$\frac{\mu_{l}(\omega=1)(1-r(1-\kappa)\hat{\zeta}(d,l))}{\mu_{l}(\omega=1)(1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=0)}\alpha(1-\kappa)\hat{\zeta}(d,l))+\mu_{l}(\omega=1)(1-r(1-\kappa)\hat{\zeta}(d,l))}$ (A.8)
It can be verified that for any $r>1$, expression A.8 is strictly larger than 0.5 if $\hat{\zeta}(d,l)=1$ (see e.g. Innocenti (2022) for a proof). We know that $\hat{\zeta}(d,l)$ goes to 1 as $d$ approaches infinity. Therefore, we can find some $\underline{d}$ large enouch such that for all larger $d$, their posterior convergse to some number larger than 0.5. Let the payoff function $v(.)$ be defined as follows, $v(x)=x$ for all $x\leq\underline{x}$ and is equal to $\underline{x}$ for all larger $x$, where $\underline{x}>\gamma_{h}$ and $\underline{x}<\gamma_{h}+\gamma_{l}\sum_{d\geq\underline{d}}p_{d}^{l}(1-\hat{\zeta}(d,l))$. For $n$ large enough, sender gets payoff $\underline{x}$ from the network strategy described above. However, with public signals, it’s impossible to always persuade some type $l$ receivers, so sender’s payoff must be less than $\underline{x}$. ∎
Proof of 2: Sender’s payoff from the strategy described in proof of Proposition 2 point 2 is at least $\frac{\overline{x}^{\beta}-1}{\beta}$, and from public signal is at most $\mu_{s}(\omega=0)(1-\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)})\frac{\gamma_{h}^{\beta}-1}{\beta}$. The former will be larger than latter for $\beta$ small enough. ∎
Proof of Proposition 3
Firstly, note that expected degree of $h$ is $E_{h}(\lambda)(\gamma_{h}E_{h}(\lambda)+\gamma_{l}E_{l}(\lambda)q)<\frac{1}{\overline{x}}$, so $E_{h}(\lambda)<\frac{1}{q\gamma_{l}E_{l}(\lambda)\overline{x}}$. This implies that given any $m>0$, the probability that a type $h$ node has connectivity smaller than $\frac{1}{q\gamma_{l}E_{l}(\lambda)\overline{x}}m$ is at least $\frac{m-1}{m}$. Select some $m$ large enough such that $1/m<\epsilon$. Also, we know that $E_{l}(\lambda)>(\overline{x}/\gamma_{l}-\frac{\gamma_{h}}{\gamma_{l}^{2}\overline{x}})^{0.5}$.
Therefore, we can find some $\overline{x}$ large enough such that expected $\lambda$ is arbitrarily large for type $l$ nodes, and at least $(1-\epsilon)$ of type $h$ nodes have connectedness weakly less than $\frac{m}{q\gamma_{l}\overline{x}E_{l}(\lambda)}$. We can look at the subnetwork formed among type $l$ nodes. As discussed above, this is equivalent to looking at a network where $\lambda$ is set to 0 for all type $h$ nodes. Therefore, by lemma 5.6 in Bollobás et al. (2007), $\rho(\infty|l,\lambda)$ is the maximal solution to the system of equations:
$\rho(\infty|l,\lambda)=1-exp(-\gamma_{l}\lambda\sum_{\lambda^{\prime}}f_{l}(\lambda^{\prime})\lambda^{\prime}\rho(\infty|l,\lambda^{\prime}))$
Fix any $x\in(0,1)$, and let $\rho(\infty|l,\lambda)=x$ for all $\lambda$. We can make $E_{l}(\lambda)$ arbitrarily large, and as it goes to infinity, the right hand sides of all equations go to 1. Therefore, LHS minus RHS is negative for all equations. If we let $\rho(\infty|l,\lambda)=1$ for all $\lambda$, the LHS minus RHS is positive for all equations. Then, by fixed point theorem (Poincare-Miranda theorem), there exists a solution to the system of equations where $\rho(\infty|l,\lambda)\geq x$ for all $\lambda$. Since $x$ is arbitrary, the size of the type $l$ giant component can be arbitrarily close to $\gamma_{l}$ for $\overline{x}$ large enough. The proportion of type $h$ nodes with degree 0 is at least $(1-\epsilon)(PO(0;\frac{\gamma_{h}}{(q\gamma_{l}E_{l}(\lambda)\overline{x})^{2}}m+\frac{m}{\overline{x}})$, which converges to $(1-\epsilon)$ as $\overline{x}$ goes to infinity for any given $m$. Therefore, we can always select $m$ and $\overline{x}$ large enough such that it’s arbitrarily close to 1.
Consider the following network signal strategy to be used for all $n$: $\pi(s|\omega=1)=1$, $\pi(s|\omega=0)=\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}$ and $\pi(\emptyset|\omega)=1-\pi(s|\omega)$. Seed is always chosen to be on $L_{1}$, the giant component. Type $h$ nodes with degree 00 always chooses action 1, and type $l$ receivers who observe signal $s$ chooses action 1. Sender’s payoff is at least
$\mu_{s}(\omega=1)v(\frac{N_{l,L_{1}}}{n}+\frac{N_{h,d=0}}{n})+\mu_{s}(\omega=0)\big{[}\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}v(\frac{N_{l,L_{1}}}{n}+\frac{N_{h,d=0}}{n})+(1-\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)})v(\frac{N_{h,d=0}}{n})\big{]}$
We have shown that, we can select $\overline{x}$ and $m$ large enough such that $\frac{N_{l,L_{1}}}{n}$ and $\frac{N_{h,d=0}}{n}$ can be arbitrarly close to $\gamma_{l}$ and $\gamma_{h}$ respectively when $n$ goes to infinity. The result follows. ∎
Proof of 3:
Point 1 is obvious from the definition of $q_{t}$. For point 2, note that degree distribution for type $t$ is a mix of poisson distributions, where the weight attached to intensity $\lambda^{\prime}(E_{t}(\lambda)\gamma_{t}+E_{t^{\prime}}(\lambda)\gamma_{t^{\prime}})$ is $f_{t}(\lambda^{\prime})$ for type $t$ receivers. Therefore, if $f_{t}(\cdot)$ FOSD $f_{t^{\prime}}(\cdot)$, then the distribution over intensities of POisson distributions for type $t$ FOSD that of type $t^{\prime}$, which in turn implies first order dominance for degree distributions.
For point 3, we use the result from Sadler (2020) that type $t$ is more connected than type $t^{\prime}$ if $\frac{\sum_{\lambda^{\prime}}f_{t}(\lambda^{\prime})PO(d;D(t,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}f_{t^{\prime}}(\lambda^{\prime})PO(d;D(t^{\prime},\lambda^{\prime}))}$ is weakly increasing in $d$. By conditions on $f_{t}$, the derivative of the ratio above has the same sign as
$\frac{\sum_{\lambda^{\prime}}f_{t^{\prime}}(\lambda^{\prime})(\lambda^{\prime})^{d}e^{-rD(t,\lambda^{\prime})}ln(rD(t,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}f_{t^{\prime}}(\lambda^{\prime})(\lambda^{\prime})^{d}e^{-rD(t,\lambda^{\prime})}}-\frac{\sum_{\lambda^{\prime}}f_{t^{\prime}}(\lambda^{\prime})(\lambda^{\prime})^{d}e^{-\lambda^{\prime}D(t^{\prime},\lambda^{\prime})}ln(D(t^{\prime},\lambda^{\prime}))}{\sum_{\lambda^{\prime}}f_{t^{\prime}}(\lambda^{\prime})(\lambda^{\prime})^{d}e^{-\lambda^{\prime}D(t^{\prime},\lambda^{\prime})}}$
Note that $ln(rD(t,\lambda))=ln(\lambda rE_{t^{\prime}}(\lambda)(\gamma_{t}r+q\gamma_{t^{\prime}}))$, and $ln(D(t^{\prime},\lambda))=ln(\lambda E_{t^{\prime}}(\lambda)(\gamma_{t}qr+\gamma_{t^{\prime}}))$. Let $\lambda_{min}$ and $\lambda_{max}$ denote the smallest and largest values that $f_{t^{\prime}}(\cdot)$ attaches positive probabilities to. The difference above is then at least $ln(rD(t,\lambda_{min}))-ln(D(t^{\prime},\lambda_{max}))$ for all $d$, which will become positive for $r$ large enough.
For the special case of the island model, the ratio above is $\big{(}\frac{\gamma_{t}+q\gamma_{t^{\prime}}}{\gamma_{t}q+\gamma_{t^{\prime}}}\big{)}^{d}\frac{e^{-\lambda^{2}(\gamma_{t}+q\gamma_{t^{\prime}})}}{e^{-\lambda^{2}(\gamma_{t}q+\gamma_{t^{\prime}})}}$, which is increasing in $d$ if $\gamma_{t}>\gamma_{t^{\prime}}$ and decreasing in $d$ if $\gamma_{t}<\gamma_{t^{\prime}}$. ∎
A.2 Proof for Voting Game
I first study the optimal public signals for the sender. If a signal doesn’t persuade type $l$ receivers, then sender gets payoff 0 when that signal is realised. Therefore, sender’s optimal strategy must maximise the probability that type $l$ receivers choose action 1. The problem is now equivalent to standard Bayesian persuasion with a single type $l$ agent. Since this is the same as standard Bayesian persuasion in Kamenica and Gentzkow (2011), the proof is omitted. Sender’s optimal payoff using public signals (in the limit as $n\rightarrow\infty$) is $V_{pub}^{*}=\mu_{s}(\omega=1)+\mu_{s}(\omega=0)\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}$
Proof of Proposition 4 :
By lemma 14, it must be the case that $\hat{\zeta}(l,d)=0$ for all $d$, or $\hat{\zeta}(l,d)>0$ for all $d>0$. Also, if it’s not the case that $\hat{\zeta}(l,d)=0$ for all $d$, then $\hat{\zeta}(l,d)$ converges to 1 as $d$ goes to infinity. This implies that as long as $\frac{\hat{L}_{1}}{n}$ doesn’t converge in probability to 0, then $d_{vote}$ is finite. One can verify that when $d_{vote}$ is finite, the ratio
$\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}\hat{\zeta}(l,d)}{1-\hat{\zeta}(l,d)}$
is weakly larger than 1 iff $d\geq d_{vote}$. By definition, the ratio is smaller than 1 for all $d$ if $d_{vote}$ is $\infty$.
I show the necessary condition first. Suppose that the necessary condition stated in proposition doesn’t hold, i.e. $\gamma_{h}+\sum_{d\geq d_{vote}}\gamma_{l}p_{d}^{l}<\overline{x}$. By definition of $d_{vote}$ and hypothesis, we can find a $\beta>0$ small enough such that
$\hat{\zeta}(l,d)+3\beta<1\quad\text{ and }\quad\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}(2\beta+\hat{\zeta}(l,d))}{1-\hat{\zeta}(l,d)-3\beta}<1$
for all $d<d_{vote}$.
We know that the proportion of type $h$ receivers converge to $\gamma_{h}$, and $\sum_{d\geq d_{vote}}\frac{N_{l,d}}{n}$ converges in probability to $\sum_{d\geq d_{vote}}\gamma_{l}p_{d}^{l}$ (by remark 1). Therefore, for fixed $\beta$ satisfying conditions above and any $\overline{d}>d_{vote}$ (if $d_{vote}$ is infinity, then select $\overline{d}$ to be large enough such that $\gamma_{h}+\sum_{d>\overline{d}}\gamma_{l}p_{d}^{l}<\overline{x}$), we can find some $n_{1}$ such that for all $n>n_{1}$, with probability at least $1-\beta$, we have $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$
Suppose that sender uses some strategy $\Pi^{n}$ in the game with $n\geq n_{1}$ receivers. Let $S_{good}$ denote the set of signals that convince both types of receivers, $S_{int}$ denote the set of signals that convince only type $h$ receivers and $S_{bad}$ as those that convince neither. Let $OB_{good}(d,\omega;\Pi^{n})$ denote the probability that a type $l$ receiver with degree $d$ observe observe signals in $S_{good}$ under $\Pi^{n}$. Similarly let $OB_{int}(d,\omega;\Pi^{n})$ denote the probability that a type $l$ receiver with degree $d$ observe observe signals in $S_{int}$ under $\Pi^{n}$. We know that the probability of observing signals in $S_{bad}$ is smaller than $\beta$. As in other proofs, let $a^{*}(\emptyset,l,d;\Pi^{n})$ denote actions of receivers with type $t$ and dgree $d$ upon observing $\emptyset$.
I show $a^{*}(\emptyset,l,d;\Pi^{n})=0$ for all $d<min\{d_{vote},\overline{d}\}$ by contradiction. Suppose that $\alpha(\emptyset,l,d;\Pi^{n})=1$ for some $d<min\{d_{vote},\overline{d}\}$. The likelihood ratio for type $l$ receiver with degree $d$ upon observing $\emptyset$ must be weakly greater than 1, therefore we must have
$\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-OB_{good}(d,\omega=1;\Pi^{n})-OB_{int}(d,\omega=1;\Pi^{n})}{1-OB_{good}(d,\omega=0;\Pi^{n})-OB_{int}(d,\omega=0;\Pi^{n})-\beta}\geq 1$
By Lemma 1, we know that $\frac{OB_{good}(d,\omega=1;\Pi^{n})}{OB_{good}(d,\omega=0;\Pi^{n})}\geq\frac{\mu_{l}(\omega=0)}{\mu_{l}(\omega=1)}$ and $\frac{OB_{int}(d,\omega=1;\Pi^{n})}{OB_{int}(d,\omega=0;\Pi^{n})}\geq\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}$. Therefore, we must have
$\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}K}{1-K-\beta}\geq 1$
for any $K\in[OB_{int}(d,\omega=0;\Pi^{n}),1-\beta)$.
$OB_{int}(d,\omega=0;\Pi^{n})$ is at most the probability of observing $s$ conditional on it being realised, which is at most $2\beta+\hat{\zeta}(l,d)$ since $g^{n}\in\mathbf{A}^{n}(\beta,\overline{d})$ with probability at least $1-\beta$. By how $\beta$ is chosen, we have $2\beta+\hat{\zeta}(l,d)<1-\beta$, which in turn implies
$\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}(2\beta+\hat{\zeta}(l,d))}{1-\hat{\zeta}(l,d)-3\beta}\geq 1$
But this inequality is a contradiction with how $\beta$ is chosen, so $a^{*}(\emptyset,l,d;\Pi^{n})=0$ for all $d<min\{d_{vote},\overline{d}\}$. Therefore, an upper bound on sender’s payoff when signal $s$ does not persuade type $l$ receivers is $v(\frac{1}{n}(N_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}N_{l,d}))$ by assuming that all type $h$ receivers take action 1, and all type $l$ receivers with degree higer than $min\{d_{vote},\overline{d}\}$ choose action 1. By remark 1, this converges in probability to $v(\gamma_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}\gamma_{l}p_{d}^{l})=0$, since $\gamma_{h}+\sum_{d\geq min\{d_{vote},\overline{d}\}}\gamma_{l}p_{d}^{l}<\overline{x}$. Therefore, sender can only get 0 unless a signal that persuades both types is realised, which implies that $\limsup_{n}V^{*}_{net}(n)$ is weakly smaller than $\sum_{\omega}\mu_{s}(\omega)\sum_{s\in S_{good}}\pi(s|\omega)$ for $n$ large enough. Since $\sum_{s\in S_{good}}\pi(s|\omega=1)\leq 1$ and
$\sum_{s\in S_{good}}\pi(s|\omega=0)\leq\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}$, this upper bound is weakly smaller than $V_{pub}^{*}$ found above. This concludes proof for the necessary condition.
To show the sufficient condition, suppose that sender uses strategy $\Pi_{2}^{n}$ described as follows for all $n$,
$\pi_{2}(s|\omega=1)=\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}\qquad\pi_{2}(s|\omega=0)=1\qquad\pi_{2}(\emptyset|\omega)=1-\pi_{2}(s|\omega)$
Let the seeds always hit $\hat{L}_{1}$. One can verify that type $h$ receivers always choose action 1, regardless of what they observe. Upon observing $\emptyset$, receivers with type $l$ and degree $d$ has likelihood ratio
$\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\pi_{2}(s|\omega=1)E(\frac{N_{\hat{L}_{1},l,d}}{N_{l,d}})}{1-\pi_{2}(s|\omega=0)E(\frac{N_{\hat{L}_{1},l,d}}{N_{l,d}})}$
By lemma 4, $\pi_{2}(s|\omega)E(\frac{N_{\hat{L}_{1},l,d}}{N_{l,d}})$ converges to $\pi_{2}(s|\omega)\hat{\zeta}(l,d)$, and thus the likelihood ratio converges to $\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}\frac{1-\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}\hat{\zeta}(l,d)}{1-\hat{\zeta}(l,d)}$. This implies that for any $d>d_{vote}$, type $l$ receivers with degree $d$ choose action 1 upon observing $\emptyset$ for $n$ large enough. Therefore, for any $\overline{d}$, we can find some $n_{2}$ such that for all larger $n$, type $l$ receivers with degrees between $d_{vote}$ and $\overline{d}$ choose action 1 upon observing $\emptyset$.
If the sufficient condition holds, we can find some $\overline{d}$ such that $\gamma_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}\gamma_{l}p_{d}^{l}(1-\hat{\zeta}(l,d))>\overline{x}$. By the argument above, for $n\geq n_{2}$, the proportion of receivers choosing action 1 in any realization is at least $\frac{1}{n}(N_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}(N_{l,d}-N_{l,d,\hat{L}_{1}}(g^{n})))$. By remark 1 and Lemma 4, it converges in probability to $\gamma_{h}+\sum_{d=d_{vote}+1}^{\overline{d}}\gamma_{l}p_{d}^{l}(1-\hat{\zeta}(l,d))>\overline{x}$, so sender’s payoff converges in probability to $v(1)$. This is strictly larger than $V_{pub}^{*}=\mu_{s}(\omega=1)v(1)+\mu_{s}(\omega=0)\frac{\mu_{l}(\omega=1)}{\mu_{l}(\omega=0)}v(1)$. ∎
A.3 Proofs for technical lemmas in section 3
Proof of Lemma 1:
Given type $t$ and degree $d$, the (subjective) probability of observing signal $s\neq\emptyset$ and being in state $\omega$ is
$\mu_{t}(\omega)\pi(s|\omega)Pr(\text{observing }s|s\text{ is realised},\omega,d,t)$
Conditional on signal $s$ and a realised network, the set of nodes being selected as seeds and thus receivers observing the signal is completely determined, and therefore does not depend on the state $\omega$. Therfore, we know that
$Pr(\text{observing }s|s\text{ is realised},\omega,d,t,g^{n})$
is independent of $\omega$. In addition, we know that conditional on $s$, the network formation process is independent of state, i.e. $Pr(\omega,g^{n},d,t|s\text{ is realised})=Pr(\omega|s\text{ is realised})Pr(g^{n},d,t|s\text{ is realised})$. This implies that $Pr(g^{n}|s\text{ is realised},\omega,d,t)$ is independent of $\omega$. Therefore, we have
$Pr(\text{observing }s|s\text{ is realised},\omega,d,t)=Pr(\text{observing }s|s\text{ is realised},d,t)$
The rest then follows from Beyesian updating. ∎
To prove the network related lemmas, I need to introduce some extra notation. I define the following Branching processes. Let $\mathbf{B}$ denote the following branching process:
  1. 1.
    The process starts with a single node in generation 1. The node has type $t$ with probability $\gamma_{t}$, and conditional on $t$ has connectivity $\lambda$ with probability $f_{t}(\lambda)$.
  2. 2.
    Given type $t$ and connectivity $\lambda$, all nodes in this process produce nodes according to Poisson distribution with intensity $D(t,\lambda)$. Recall that $D(t,\lambda)\equiv\lambda(\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda))$.
  3. 3.
    Conditional on type $t$, connectivity $\lambda$ and having $d$ offsprings, the distribution over types and connectivity of these offsprings is given by a multinomial distribution with $d$ trials, and success probability for pair $(t^{\prime},\lambda^{\prime})$ is $\frac{\lambda\lambda^{\prime}f_{t^{\prime}}(\lambda^{\prime})\gamma_{t^{\prime}}}{D(t,\lambda)}$ if $t=t^{\prime}$, and $\frac{\lambda\lambda^{\prime}f_{t^{\prime}}(\lambda^{\prime})\gamma_{t^{\prime}}q}{D(t,\lambda)}$ if $t\neq t^{\prime}$.
  4. 4.
    All nodes produce offsprings mutually independently of each other.
Also, let $\rho(<m)$ and $\rho(m)$ denote the probability that the process contains less than $m$ and exactly $m$ nodes. Let $\rho(\infty)$ denote the probability that the process never dies out. One can show that $\rho(\infty)=1-\lim_{m\rightarrow\infty}\rho(<m)$, see e.g. Bollobás et al. (2007). Similarly, let $\rho(<m|\cdot)$, $\rho(m|\cdot)$ and $\rho(\infty|\cdot)$ denote the respective probabilities conditional on properties of the initial node.
Let $N_{<m}(g^{n})$ denote the number of nodes on components of size strictly less than $m$ in the realised network $g^{n}$. Let $N_{L_{1}}(g^{n})$ denote the number of nodes on $L_{1}$.
I’ll prove lemmas Lemma 3 and Lemma 2 together in the following proof.
Proof of Lemma 3 and Lemma 2
An upper bound on the proportion of receivers outside $L_{1}$ that observe $s$ can be obtained by assuming that all receivers on components larger than size $\overline{m}$ observes $s$, and all seeds are on separate components of size $\overline{m}$. Summing these up, the proportion of receivers not on $L_{1}$ who can observe $s$ is at most
$\frac{n-N_{<\overline{m}}(g^{n})-N_{L_{1}}(g^{n})}{n}+\frac{n^{\alpha}\overline{m}}{n}$
We know that $N_{<\overline{m}}(g^{n})/n$ converges in probability to (by Theorem 9.1 from Bollobás et al. (2007)) $\rho(<\overline{m})$, and $\frac{N_{L_{1}}}{n}$ converges in probability to $\rho(\infty)$. For any $\epsilon>0$, We can select $\overline{m}$ large enough such that $1-\rho(<\overline{m})-\rho(\infty)$ is smaller than $\epsilon$. For any given $\overline{m}$, $\frac{n^{\alpha}\overline{m}}{n}$ converges to 0. The first parts of Lemma 3 and Lemma 2 then follow since $\epsilon$ can be arbitrary.
Next, we consider signals that persuade only type $h$ receivers and thus only passed on by them. We first bound the proportion of type $h$ nodes not on $\hat{L}_{1}$ that can observe this signal. By the argument preceeding 1, to study the type $h$ subnetwork, we can essentially study the transformed network formation process where $\lambda_{i}=0$ for all type $l$ receivers. Then, applying the same steps from last paragraph shows the proportion of type $h$ nodes that are not on $\hat{L}_{1}$ and observe $s$ converges in probability to 0, if $s$ persuades only type $h$ receivers.
A type $l$ node can receive the signal only if they have some type $h$ neighbour who receives the signal. Therefore, we proceed by bounding the number of neighbours of type $h$ nodes who receive signal and are not on $\hat{L}_{1}$. Since we know that $\sum_{d}p_{d}^{h}d$ is bounded, we can find some $\overline{d}>0$ such that $\sum_{d\geq\overline{d}}p_{d}^{h}d<\epsilon$.
The proportion of type $h$ receivers who are not on $\hat{L}_{1}$ and observe $s$ converges in probability to 0, so the proportion of type $h$ receivers that in addition has less than $\overline{d}$ neighbors must also converge to 0. This in turn implies that the number of neighbours they have converge in probability to 0. Therefore, we can provide the upper bound by assuming that all type $h$ nodes with degree higher than $\overline{d}$ observe the signal and are not on $\hat{L}_{1}$ [^12] For a signal $s$ that only persuades type $h$ receivers, if no seeds are on $\hat{L}_{1}$ or a type $l$ node has no neighbour on $\hat{L}_{1}$, then the type $l$ node can only get the signal $s$ from type $h$ receivers who are not on $\hat{L}_{1}$. The number of neighbors they have is $\sum_{d\geq\overline{d}}N_{h,d}d$, and the ratio $\sum_{d\geq\overline{d}}\frac{N_{h,d}d}{n}$ converges in probability to $\sum_{d\geq\overline{d}}p_{d}^{h}d<\epsilon$. Since $\epsilon$ is arbitrary, the second parts of Lemma 3 and Lemma 2 follow. ∎
To prove lemma 4, I first prove the following lemma, which says that if we look at local properties, the random network is approximated by the branching process above. Let $N_{t,d,\lambda,m}(g^{n})$ denote the number of nodes with type $t$, connectedness $\lambda$, degree $d$ and is on a component of size $m$ in network $g^{n}$.
Lemma 9.
$\frac{N_{t,d,\lambda,m}(G^{n})}{n}$ converges in probability to $\gamma_{t}f_{t}(\lambda)Poisson(d;D(t,\lambda))\rho(m|t,d,\lambda)$.
The proof of this largely follows from Sadler (2020)’s proof for Theorem 10. Let $g_{i,r}^{n}$ denote the rooted graph formed by deleting all nodes that are more than $r$ steps away from $i$, and letting $i$ be the root. Note that given any realised branching process, we can connect each node to their offsprings and parent. Then, we can form a rooted graph and delete all nodes that are distance $r$ away from the initial node, and let the initial node be the root.
For these proofs, I introduce the following notation. Let $\eta(t,t^{\prime},\lambda,\lambda^{\prime})$ be $\lambda\lambda^{\prime}f_{t^{\prime}}(\lambda^{\prime})\gamma_{t^{\prime}}$ if $t=t^{\prime}$ and $q\lambda\lambda^{\prime}f_{t^{\prime}}(\lambda^{\prime})\gamma_{t^{\prime}}$ if $t\neq t^{\prime}$. This $\eta(\cdot)$ function captures the expected number of a node’s neighbours of type $t^{\prime}$ and connectedness $\lambda^{\prime}$, given the node’s type $t$ and connectedness $\lambda$.
Lemma 10.
Let $\mathbf{P}^{r}$ be a finite set of rooted graph where all nodes are within distance $r$ of the root. Uniformly randomly select some node $i$, the probability that $g_{i,r}^{n}\in\mathbf{P}^{r}$ converges to the probability that branching process $B$ produces a rooted graph in $\mathbf{P}^{r}$.
Proof.
It’s sufficient to show that the claim is true for all $\mathbf{P}^{r}$ that’s a singleton, i.e. it contains only one rooted graph.
I do this by induction. Suppose that the claim above is true for $r$, I show that it must be true for $r+1$. Let $g_{r+1}$ be any rooted graph where all nodes are within distance $r+1$ of the root, and there’s at least 1 node with distance $r+1$ from the root. Let $g_{r}$ denote the graph after deleting nodes with distance $r+1$ from the root node. Given the induction hypothesis, it’s sufficient to show that $Pr(g_{i,r+1}^{n}=g_{r+1}|g_{i,r}^{n}=g_{r})$ converges to the probability that the branching process produces $g_{r+1}$ conditional on producing $g_{r}$.
Let $I$ be the set of nodes with distance $r$ from the root node in $g_{r}$. Let $d_{j,t^{\prime},\lambda^{\prime}}$ denote the number of neighbours of node $j$ in $g_{r+1}$ that have type $t^{\prime}$, connectedness $\lambda^{\prime}$ and are at distance $r+1$ from root. Note that by definition, $g_{r+1}$ is fully described by $g_{r}$ and $\{(d_{j,t^{\prime},\lambda^{\prime}})_{t^{\prime},\lambda^{\prime}}|j\in I\}$.
First, we look at the probability that the branching process produces $\{(d_{j,t^{\prime},\lambda^{\prime}})_{t^{\prime},\lambda^{\prime}}|j\in I\}$ conditional on $g_{r}$. Conditional on type $t_{j}$ and connectedness $\lambda_{j}$ of any node $j\in I$, the probability that it produces $(d_{j,t^{\prime},\lambda^{\prime}})_{t^{\prime},\lambda^{\prime}}$ is $\Pi_{t^{\prime},\lambda^{\prime}}PO(d_{j,t^{\prime},\lambda^{\prime}};\eta(t_{j},t^{\prime},\lambda_{j},\lambda^{\prime}))$, and it’s independent of $g_{r}$ and how other nodes in $I$ produce nodes. Therefore, the probability that the branching process produces $\{(d_{j,t^{\prime},\lambda^{\prime}})_{t^{\prime},\lambda^{\prime}}|j\in I\}$, and thus $g_{r+1}$, conditional on $g_{r}$ is
$\Pi_{j\in I}\Pi_{t^{\prime},\lambda^{\prime}}PO(d_{j,t^{\prime},\lambda^{\prime}};\eta(t_{j},t^{\prime},\lambda_{j},\lambda^{\prime}))$ (A.9)
Now we consider the network. Conditional on types and connectedness of all nodes, connection between any pair is independent of connections between other pairs. Let $N_{t^{\prime},\lambda^{\prime}}$ denote the number of nodes not in $I$ that have type $t^{\prime}$ and connectedness $\lambda^{\prime}$. Also, note that neighbours of nodes in $I$ that are at distance $r+1$ from root must be nodes not in $g_{r}$. Therefore, conditional on types and connectedness of all nodes and $g_{r}$, the probability of producing $\{(d_{j,t^{\prime},\lambda^{\prime}})_{t^{\prime},\lambda^{\prime}}|j\in I\}$ is
$\Pi_{j\in I}\Pi_{t^{\prime},\lambda^{\prime}}Binom(d_{j,t^{\prime},\lambda^{\prime}}|N_{t^{\prime},\lambda^{\prime}},\frac{\mathbf{1}(t_{j}=t^{\prime})+\mathbf{1}(t_{j}\neq t^{\prime})q}{n}\lambda_{j}\lambda^{\prime})$
Since $N_{t^{\prime},\lambda^{\prime}}\frac{\mathbf{1}(t_{j}=t^{\prime})+\mathbf{1}(t_{j}\neq t^{\prime})q}{n}\lambda_{j}\lambda^{\prime}$ converges in probability to $\eta(t_{j},t^{\prime},\lambda_{j},\lambda^{\prime})$, so the expression above converges in probability to expression A.9.
To complete the induction argument, we need to show that the claim holds true for $r=0$. For $r=0$, $g_{r}$ consists only of the root node, and thus is fully described by the type and connectedness of the node. The probability that the branching process produces a node with type $t$ and connectedness $\lambda$ is $\gamma_{t}f_{t}(\lambda)$. From our network formation process, as $n$ gets large, the probability that a randomly selected node has type $t$ and connectedness $\lambda$ has the same expression. ∎
This shows that the expected proportion of nodes with any given local propoerty converges to the probabilty that $\mathbf{B}$ has this property. We wish to show that they converge in probability. Note that the proportion of nodes with type $t$ and connectedness $\lambda$ converges in probability to $\gamma_{t}f_{t}(\lambda)$. By remark 8.8 and also lemma A.5 in Bollobás et al. (2007), we can assume that connectedness for receivers are deterministic, and the proportion of nodes with type $t$ and connectedness $\lambda$ converges to $\gamma_{t}f_{t}(\lambda)$. This will not change the asymptotic distribution over networks. I will use this assumption for the rest of the proof for Lemma 9
For any 2 realised networks $g_{1}^{n}$ and $g_{2}^{n}$ of the same size, I say that they are related by a deletion if one of them can be obtained by deleting one link from the other. Let $F(\cdot)$ be a real-valued function defined on the possible networks . We say that $F(\cdot)$ is $\beta$-Lipschitz if for any two networks related by a deletion, we have that $|F(g_{1}^{n})-F(g_{2}^{n})|<\beta$.
Lemma 11.
If $F$ is a $\beta$-Lipschitz function, then $F(g^{n})$ converges in probability to its expected value. Formally, for any $\epsilon>0$,
$Pr(|F(g^{n})-E(F(g^{n}))|\geq\epsilon)\leq 2e^{-\frac{\epsilon^{2}}{2\beta^{2}n}}$
Proof.
We consider a random process in which we seqentially reveal whether two nodes are connected. For each agent $i$, we sequentially reveal whether $i$ is connected to nodes with larger index. Note that starting from $i=1$ and going to $i=n$ reveals the whole network.
Let $M_{j}$ denote the realised connection patterns generated by the revelation process upto step $j$. The process $X_{j}=E(F(g^{n})|M_{j})$ is a martingale [^13] To see this, note that it’s sufficient to show that $E(E(F(g^{n})|M_{j+1}|M_{j})=E(F(g^{n})|M_{j})$, which holds almost by definition. Also, we know that $|E(F(g^{n})|M_{j})-E(F(g^{n})|M_{j+1})|<\beta$. This holds since revelation of connections for steps later than $j+1$ are independent of realizations before. Therefore, given any $M_{j}$, the distributions over connections revealed after $j+1$-th step is the same as conditioning on $M_{j+1}$. By construction of the process, $M_{j}$ and $M_{j+1}$ can only differ at most by deletion of one link. This inequality along with Azuma’s inequality produces the result. ∎
To show Lemma 9, we need to show that $N_{t,d,\lambda,m}(g^{n})$ are Lipschitz, which is shown in the following lemma.
Lemma 12.
Let $N_{t,d,\lambda,m}(g^{n})$ is $22m+$ 2-Lipschitz.
Proof.
Consider any network $g^{n}$ and delete some link. Note that if removing the link doesn’t change the component structure. Then it can only change the degree of two agents and therefore will only change the number of nodes with the described property by 2.
If it changes the component structure. It can only break some component into 2. If that component that’s broken is of size $m$, then it will change the number of nodes with the property described by at most $m+2$. If it’s of size more than $m$, it’s possible that it’s broken into two compoenents of size $m$. Therefore, it can change the number of nodes with the desired property by at most $22m+$ 2. ∎
Combining Lemma 10, Lemma 11 and Lemma 12 gives Lemma 9. Using Lemma 9, we proof Lemma 4 in the following two lemmas, also using the following remark.
Remark 4.
$\rho(\infty|t,d,\lambda)$ is given by
$1-(\frac{\sum_{\lambda^{\prime}}(1-\rho(\infty|t,\lambda^{\prime}))\gamma_{t}f_{t}(\lambda^{\prime})\lambda^{\prime}+(1-\rho(\infty|t^{\prime},\lambda^{\prime}))q\gamma_{t^{\prime}}f_{t^{\prime}}(\lambda^{\prime})\lambda^{\prime}}{\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda)})^{d}$
Note that this expression does not depend on $\lambda$, so $\rho(\infty|t,d)$ is equal to the same expression.
Proof.
This is standard in the study of branching processes (see. e.g. sections 5 and 6 in Bollobás et al. (2007)) and I provide a brief argument here. The process never dies out if and only if at least one process that continues from one of the second generation offsprings does not die out.
Conditional on the initial node having $(t,\lambda)$, each second generation node has type $(t,\lambda^{\prime})$ with probability $\frac{\gamma_{t}f_{t}(\lambda^{\prime})\lambda^{\prime}}{\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda)}$ and $(t^{\prime},\lambda^{\prime})$ with probability $\frac{\gamma_{t^{\prime}}f_{t^{\prime}}(\lambda^{\prime})q\lambda^{\prime}}{\gamma_{t}E_{t}(\lambda)+q\gamma_{t^{\prime}}E_{t^{\prime}}(\lambda)}$, independent of types and connectedness of other nodes. From construction, we know that if a second generation node has type $(t,\lambda)$, then the process that continues from it behaves in the same way as $\mathbf{B}$ with an initial node of type $t$ and connectedness $\lambda$. Therefore, each second generation node with type $t$ and connectedness $\lambda$ has a probability of dying out $1-\rho(\infty|t,\lambda)$. Since all nodes produce offsprings mutually independently from each other, conditional on types and connectedness of the second generation nodes, the processes that continue from them are also mutually independent. Then, the expression follows. ∎
Lemma 13.
Conditional on having degree $d$ and type $t$ and $\lambda$ with $f_{t}(\lambda)>0$, the proportion of nodes on the $L_{1}$ converges in probability to $\rho(\infty|t,d)$, which is increasing in $d$.
Proof.
Let $N_{t,d,\lambda,\leq m}(g^{n})$ be the number of nodes that have type $t$, degree $d$, connectedness $\lambda$ and is on a component of size smaller than $m$. Let $N_{t,d,\lambda}(g^{n})$ be the number of nodes that have type $t$, degree $d$ and connectedness $\lambda$.
By Lemma 9, $N_{t,d,\lambda,\leq m}(g^{n})/N_{t,d,\lambda}(g^{n})$ converges in probability to $\rho_{D}(\leq m|t,d,\lambda)$. Let $N_{t,d,\lambda,L_{1}}(g^{n})$ denote the number of nodes that have type $t$, degree $d$, connectedness $\lambda$ and are on the largest component $L_{1}$.
By Theorem 9.1 from Bollobás et al. (2007), $\frac{N_{>m}(g^{n})-N_{L_{1}}(g^{n})}{N_{t,d,\lambda}}$ converges to in probability to $\frac{1-\rho(\leq m)-\rho(\infty)}{\gamma_{t}f_{t}(\lambda)PO(d;D(t,\lambda))}$. For any $\epsilon>0$, we can find some $\overline{m}$ large enough such that $\frac{1-\rho(\leq m)-\rho(\infty)}{\gamma_{t}f_{t}(\lambda)PO(d;D(t,\lambda))}<\epsilon$ for all $m\geq\overline{m}$. We can also fine $\overline{m}$ large enough such that $1-\rho_{D}(\leq\overline{m}|t,\lambda,d)\leq\rho_{D}(\infty|t,\lambda,d)+\epsilon$. By definition, we know that
$\frac{N_{t,d,\lambda,L_{1}}(g^{n})}{N_{t,d,\lambda}}+\frac{N_{t,d,\lambda,\leq\overline{m}}(g^{n})}{N_{t,d,\lambda}}+\frac{-N_{t,d,\lambda,L_{1}}(g^{n})+N_{t,d,\lambda,>\overline{m}}(g^{n})}{N_{t,d,\lambda}}=1$
The probability that $\frac{N_{t,d,\lambda,\leq\overline{m}}(g^{n})}{N_{t,d,\lambda}}$ is smaller than $\rho_{D}(\leq\overline{m}|t,d,\lambda)-\epsilon$ converges to 0. Therefore, the probability that
$\frac{N_{t,d,\lambda,L_{1}}(g^{n})}{N_{t,d,\lambda}}\geq\rho_{D}(\infty|t,d,\lambda)+2\epsilon\geq 1-\rho_{D}(\leq\overline{m}|t,d,\lambda)+\epsilon$
converges to 0. Also, since the probabilities that $\frac{N_{t,d,\lambda,\leq\overline{m}}(g^{n})}{N_{t,d,\lambda}}>\rho_{D}(\leq\overline{m}|t,d,\lambda)+\epsilon$ and $\frac{-N_{t,d,\lambda,L_{1}}(g^{n})+N_{t,d,\lambda,>\overline{m}}(g^{n})}{N_{t,d,\lambda}}>\epsilon$ converge to 0. Therefore, the probability that
$\frac{N_{t,d,\lambda,L_{1}}(g^{n})}{N_{t,d,\lambda}}\leq 1-\rho_{D}(\leq\overline{m}|t,d,\lambda)-2\epsilon$
converges to 0. Therefore, the probability that $\frac{N_{t,d,\lambda,L_{1}}(g^{n})}{N_{t,d,\lambda}}\leq\rho_{D}(\infty|t,d,\lambda)-2\epsilon$ converges to 0. Since $\epsilon$ is arbitrary, this shows that $\frac{N_{t,d,\lambda,L_{1}}(g^{n})}{N_{t,d,\lambda}}$ converges in probability to $\rho_{D}(\infty|t,d,\lambda)$. This also implies that $\frac{N_{t,d,L_{1}}(g^{n})}{N_{t,d}}$ converges in probability to $\rho_{D}(\infty|t,d)$. The fact that $\rho_{D}(\infty|t,d)$ is increasing in $d$ then follows from remark 4. ∎
Lemma 14.
Conditional on having degree $d$ and type $t$, the proportion of nodes with a neighbour on the $\hat{L}_{1}$ converges in probability, and the limit is increasing in $d$
Proof.
Using the argument preceeding 1 again, the distribution over subnetwork among type $h$ nodes is the same as the distribution over networks after setting $\lambda_{i}=0$ for all type $l$ nodes. I use $\hat{\rho}(\infty)$ to denote the probability that the associated process never dies out, and similarly $\hat{\rho}(\infty|\cdot)$ to denote the conditional probabilities.
Let $N_{h,d_{h},\lambda}$ denote the number of nodes with type $h$, $d_{h}$ type $h$ neighbours and connectedness $\lambda$, and let $N_{h,d_{h},\lambda,\hat{L}_{1}}$ denote those that in addition are on $\hat{L}_{1}$. Using Lemma 13, we get that $\frac{N_{h,d_{h},\lambda,\hat{L}_{1}}(g^{n})}{N_{h,\lambda,d_{h}}}$ converges in probability to $\hat{\rho}(\infty|h,d_{h},\lambda)$.
A node of type $h$ and connectedness $\lambda$ is connected to some randomly selected type $l$ node with probability $\frac{\lambda E_{l}(\lambda)q}{n}$. Also, whether it’s connected is independent of all the other connections. This implies that the probability that it has $d_{l}$ type $l$ neighbours is $Binom(d_{l};N_{l},\frac{\lambda\sum_{\lambda^{\prime}}\lambda^{\prime}f_{l}(\lambda^{\prime})q}{n})$, where $N_{l}$ is the number of type $l$ receivers. Therefore, the number of nodes with type $h$, connectedness $\lambda$, $d_{h}$ high type neighbours, $d_{l}$ type $l$ neighbours and on $\hat{L}_{1}$ can be seen as the random variable formed in the following way: We first form the type $h$ subnetwork as described above, and then take $N_{h,d_{h},\lambda,\hat{L}_{1}}$ independent draws of the binomial random variable with $N_{l}$ trials and success probability $\frac{\lambda E_{l}(\lambda)q}{n}$.
We know that $Binom(d_{l};N_{l},\frac{\lambda E_{l}(\lambda)q}{n})$ converges to $PO(d_{l};\gamma_{l}\lambda E_{l}(\lambda)q)$, and $N_{h,d_{h},\lambda,\hat{L}_{1}}$ goes to infinity as $n$ goes to infinity, therefore
$\frac{N_{h,d_{h},d_{l},\lambda,\hat{L}_{1}}}{n}\rightarrow_{p}\gamma_{h}f_{h}(\lambda)PO(d_{h};\lambda\gamma_{h}E_{h}(\lambda))\hat{\rho}(\infty|d_{h},h,\lambda)PO(d_{l},\lambda q\gamma_{l}E_{l}(\lambda))$
Similarly, we can use this logic to get $N_{h,\lambda,d_{h},d_{l}}$, i.e., We first form the type $h$ subnetwork as described above, and then take $N_{h,d_{h},\lambda}$ independent draws of the binomial random variable with $N_{l}$ trials and success probability $\frac{\lambda E_{l}(\lambda)q}{n}$. We have
$\frac{N_{h,\lambda,d_{h},d_{l}}}{n}\rightarrow_{p}\gamma_{h}f_{h}(\lambda)PO(d_{h};\lambda\gamma_{h}E_{h}(\lambda))PO(d_{l},\lambda q\gamma_{l}E_{l}(\lambda))$
From these, we have $\frac{N_{h,\lambda,d_{h},d_{l},\hat{L}_{1}}}{N_{h,\lambda,d_{h},d_{l}}}\rightarrow_{p}\hat{\rho}(\infty|d_{h},h,\lambda)=\hat{\rho}(\infty|d_{h},h)$, which also implies that $\frac{N_{h,d_{h},d_{l},\hat{L}_{1}}}{N_{h,d_{h},d_{l}}}$ converges in probability to $\hat{\rho}(\infty|d_{h},h)$. Since $\frac{N_{h,d_{l},d_{h}}}{N_{h,d}}$ converges in probability to $Binom(d_{h};d,q_{h})$, we know that
$\frac{N_{h,d,\hat{L}_{1}}(g^{n})}{N_{h,d}(g^{n})}\rightarrow_{p}\sum_{d_{h}}Binom(d_{h};d,q_{h})\hat{\rho}(\infty|d_{h},h)$ (A.10)
This concludes the proof for type $h$ nodes and we now consider type $l$ nodes. Let $N_{h,\lambda^{\prime},\hat{L}_{1}}(g^{n})$ denote nodes with type $h$, connectedness $\lambda^{\prime}$ and are on $\hat{L}_{1}$, and let $N_{h,\lambda^{\prime}}(g^{n})$ denote nodes with type $h$, connectedness $\lambda^{\prime}$.
The probability that a node of type $l$ and connectedness $\lambda$ is connected to a node of type $h$ and connectedness $\lambda^{\prime}$ is $\frac{\lambda\lambda^{\prime}q}{n}$. Given types and connectedness of a pair of nodes, whether they are connected are independent of the types of all other nodes, and formation of all other links. Therefore, conditional on the type $h$ subnetwork, a node of type $l$, connectedness $\lambda$ and $d_{l}$ type $l$ neighbours is connected to 00 nodes on $\hat{L}_{1}$ and $k$ type $h$ nodes not on $\hat{L}_{1}$ with probability
$\Pi_{\lambda^{\prime}}Binom(0;N_{h,\lambda^{\prime},\hat{L}_{1}},\frac{\lambda\lambda^{\prime}q}{n})\sum_{\sum_{\lambda^{\prime}}k_{\lambda^{\prime}}=k}\Pi_{\lambda^{\prime}}Binom(k_{\lambda^{\prime}};N_{h,\lambda^{\prime}}-N_{h,\lambda^{\prime},\hat{L}_{1}},\frac{q\lambda\lambda^{\prime}}{n})$ (A.11)
Since $\frac{N_{h,\lambda^{\prime},\hat{L}_{1}}}{n}$ and $\frac{N_{h,\lambda^{\prime}}}{n}$ converges in probability to $\gamma_{h}f_{h}(\lambda^{\prime})\hat{\rho}(\infty|h,\lambda^{\prime})$ and $\gamma_{h}f_{h}(\lambda^{\prime})$ respectively, using properties of Poisson distribution, we know that A.11 converges to
$PO(k;\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))Binom(k;k,\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})(1-\hat{\rho}(\infty|h,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})})$
Using this, and the fact that $N_{l,\lambda,d_{l}}$ goes to infinity as $n$ goes to infinity, we have
$\frac{N_{l,d_{l},d_{h},\lambda}-N_{l,d_{l},d_{h},\lambda,\hat{L}_{1}}}{N_{l,d_{l},\lambda}}=PO(d_{h};\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))Binom(d_{h};d_{h},\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})(1-\hat{\rho}(\infty|h,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})})$
Similarly, conditional on the type $h$ subnetwork, a node of type $l$, connectedness $\lambda$ and $d_{l}$ type $l$ neighbours is connected to $k$ type $h$ nodes with probability
$\sum_{\sum_{\lambda^{\prime}}k_{\lambda^{\prime}}=k}\Pi_{\lambda^{\prime}}Binom(k_{\lambda^{\prime}};N_{h,\lambda^{\prime}},\frac{q\lambda\lambda^{\prime}}{n})$
which converges to $PO(k;\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))$. Therefore, $\frac{N_{l,d_{l},d_{h},\lambda}}{N_{l,d_{l},\lambda}}$ converges in probability to $PO(d_{h};\sum_{\lambda^{\prime}}\lambda\lambda^{\prime}q\gamma_{h}f_{h}(\lambda^{\prime}))$. Therefore,
$\frac{N_{l,d_{l},d_{h},\lambda,\hat{L}_{1}}}{N_{l,d_{l},d_{h},\lambda}}=1-(\frac{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})(1-\hat{\rho}(\infty|h,\lambda^{\prime}))}{\sum_{\lambda^{\prime}}\lambda^{\prime}f_{h}(\lambda^{\prime})})^{d_{h}}=\hat{\rho}(\infty|d_{h},h)$
where the last equality follows from 4 and constuction of the network formatio process associated with $\hat{\rho}(\cdot)$. This implies that $\frac{N_{l,d_{l},d_{h},\hat{L}_{1}}}{N_{l,d_{l},d_{h}}}=\hat{\rho}(\infty|d_{h},h)$, and
$\frac{N_{l,d,\hat{L}_{1}}}{N_{l,d}}\rightarrow_{p}\sum_{d_{h}}Binom(d_{h};d,1-q_{l})\hat{\rho}(\infty|d_{h},h)$
The fact that both these limits are increasing follows since $\rho(\infty|d_{h},h)$ is increasing in $d_{h}$, and binomial distribution with a larger number of trials FOSD one with a smaller number of trials, given the same success probability. ∎